### v1

Abstract
This paper examines the identification and estimation of treatment effects in staggered adoption designs – a common extension of the canonical Difference-in-Differences (DiD) model to multiple groups and time-periods – in the presence of (time varying) misclassification of the treatment status as well as of anticipation. We demonstrate that standard estimators are biased with respect to commonly used causal parameters of interest under such forms of misspecification. To address this issue, we provide modified estimators that recover the Average Treatment Effect of observed and true switching units, respectively. Additionally, we suggest a testing procedure aimed at detecting the timing and extent of misclassification and anticipation effects. We illustrate the proposed methods with an application to the effects of an anti-cheating policy on school mean test scores in high stakes national exams in Indonesia.
JEL Classification: C13, C21, C51
Keywords: Measurement Error, Causal Inference, Moment Equality Tests, Misspecification Tests
1 Introduction
The Difference-in-Differences (DiD) methodology is arguably one of the most popular research designs in applied economics, offering a simple and intuitive framework to estimate causal effects in observational settings. It serves as a preferred tool for policy evaluations, labor market studies, and other areas where randomized controlled trials are impractical. Staggered adoption designs naturally extend the canonical two-by-two DiD framework to settings where multiple groups adopt treatments at different times. This structure enables researchers to exploit variation in treatment timing and group-level adoption, aggregate treatment effects across groups, and accommodate heterogeneity in treatment effects. [^1] To illustrate the prevalence of staggered adoption designs, Baker et al. (2022) report that more than 54% of all DiD studies published in the leading Finance and Accounting journals between 2000 and 2019 (744 studies in total) were based on staggered adoption designs.
Despite the widespread use of staggered adoption designs in practice, common forms of misspecification, such as (time-varying) treatment status misclassification or anticipatory effects, have received relatively little attention, even though they can undermine the validity of estimation and inference procedures in this setting. In this paper, we address these challenges and examine the implications of time-varying misclassification and anticipation for the identification and estimation of commonly used estimands in staggered adoption designs. In a second step, we propose remedies by introducing modified estimators that can recover well-defined treatment parameters, such as the Average Treatment Effect (ATE) of units in the observed switching-to-treatment cells, or of units in the actual or true switching-to-treatment cells. The former parameter measures the ATE for all treated units (switcher-to-treatment and already treated units) at the time when treatment is observed in the data, while the latter is the ATE of all switching-to-treatment units at the time when they actually switch status, which may be unobserved in the data. Additionally, we propose a testing procedure aimed at detecting the timing and extent of misclassification or anticipation effects.
Misclassification of the treatment status arises when the recorded treatment status does not coincide with the actual treatment status. This can occur due to data limitations, such as when the exact timing of treatment is not observed (e.g., Bindler and Hjalmarsson, 2018), or due to reporting and classification errors when treatment status can only be inferred indirectly (see Goldschmidt and Schmieder, 2017). Anticipation effects, on the other hand, occur when individuals or units are either exposed to treatment prior to the recorded treatment date or adjust their behavior in anticipation of future treatment, thereby contaminating pre-treatment periods (see, e.g., Barrios, 2022; Bismarck-Osten et al., 2022; Berkhout et al., 2024). Both issues are particularly problematic in staggered adoption designs, where treatment assignment evolves over time and across groups, amplifying the risk of “forbidden comparisons” and leading to invalid inference, even when estimands appropriate for staggered adoption settings are targeted.
This paper makes several contributions to the literature. First, it characterizes the bias introduced by misclassification and anticipation effects in commonly used estimators across different settings. Specifically, under the restrictive assumption of homogeneous treatment effects, we show that the Two-Way Fixed Effects (TWFE) estimator, the most prominent estimator in standard two-by-two DiD settings, is biased. The bias arises because misspecification distorts the observed timing and incidence of treatment, leading to “forbidden comparisons” where treated units may be incorrectly classified or compared to units that have already been partially treated. We then extend our analysis to alternative DiD estimands designed to accommodate heterogeneous treatment effects in staggered adoption designs. In particular, we consider the ATE of observed switching-to-treatment cells as well as the ATE of actual switching-to-treatment cells. For both causal parameters, we derive the structure and nature of the bias under staggered adoption, highlighting the limitations of existing methods in the presence of such forms of misspecification.
As a second contribution, we define alternative estimators that allow to recover the aforementioned causal parameters in staggered adoption designs. In particular, we introduce a bias-corrected version of the staggered adoption estimator proposed by De Chaisemartin and D’Haultefoeuille (2020), which recovers the ATE of observed switching-to-treatment units. Based on knowledge of the timing and extent of misspecification, the estimator adjusts the comparisons conducted by the original estimator of De Chaisemartin and D’Haultefoeuille (2020) by using groups unaffected by misspecification. In contrast, for the ATE of actual switching-to-treatment cells, additional steps are required. Since the actual timing of treatment adoption is unobserved at the unit level, we exploit a homogeneity assumption on the misclassification (anticipation) probability, which restricts heterogeneity across units but not over time. This allows us to consistently estimate the fraction of actual switching-to-treatment units by extending the Parallel Trends (PT) assumption to the (potential) outcomes under treatment. In a subsequent step, we construct an estimator for the ATE of actual switching-to-treatment cells and derive its asymptotic properties.
As a third contribution, we propose a two-step testing procedure to detect the presence, extent, and timing of misspecification. These tests provide guidance on whether and when misclassification or anticipation occurs, which is required by the estimators introduced in the previous section. In particular, the corrections proposed in this paper rely on common PT assumptions. However, distinguishing between bias arising from misspecification and bias due to PT violations in pre-treatment periods is inherently challenging. To address this, we propose a two-step moment equality testing framework that combines a pre-trend test with a test for misspecification. The first step aims to maximize confidence in the PT assumption by examining pre-treatment evidence for its validity, comparing not-yet-switched comparison groups with never treated groups. Conditional on passing this stage, the second step tests for misclassification or anticipation by comparing observed switcher groups to not-yet-switched groups using lagged periods. We implement these tests as moment equality tests and derive their asymptotic properties.
To illustrate the finite sample performance of both the estimators and the tests, we carry out a small-scale Monte Carlo simulation. The findings suggest a good performance even in cases of misspecification because of misclassification or anticipation. We then apply these methods to assess the effects and the presence of anticipation in a study by Berkhout et al. (2024) on the effects of a new computer based testing system on school mean test scores in Indonesia between 2015 to 2019. The testing system, which was voluntary and intended to curb cheating behavior, was reported to have led schools to anticipate measures that limit cheating. Our estimates for the ATE of observed switching-to-treatment units confirm this and align with Berkhout et al. (2024), while the estimates for the ATE of actual switching-to-treatment units are slightly larger in magnitude. Moreover, we estimate the fraction of anticipating schools to be around 15% in 2018, and up to 30% in 2019. Finally, the misspecification tests provide evidence for anticipation effects, though they also suggest some violation of parallel pre-trends in periods just prior to that anticipation for at least one cohort.
By providing a framework for inference in staggered adoption designs under misspecification, this paper contributes to the growing literature on the methodological underpinnings of DiD. In particular, the paper builds on three strands of the econometric literature. First, it contributes to the growing body of research on DiD methods tailored to staggered adoption designs, including recent advancements that address the challenges of heterogeneous treatment effects and dynamic treatment timing (e.g., De Chaisemartin and D’Haultefoeuille, 2020; Callaway and Sant’Anna, 2021; Borusyak et al., 2024). Second, the paper intersects with a broader literature on the misclassification of treatment status, which is well known to compromise the validity of inference procedures (e.g., Lewbel, 2007b; Ura, 2018; Denteh and Kedagni, 2022). Unlike other approaches in the DiD literature that rely on proxy variables (e.g., Botosaru and Gutierrez, 2018) or exclusion restrictions (Negi and Negi, 2025), our framework does not require such external information. Third, the paper relates to recent contributions on pre-testing of Parallel Trends and its implications for causal inference depending on the test outcome (e.g., Rambachan and Roth, 2023; De Chaisemartin and D’Haultefoeuille, 2024).
The paper is structured as follows. Section 2 introduces the outcome framework together with the misclassification design and outlines the key assumptions. Section 3 analyzes the bias of commonly defined estimands under homogeneous and heterogeneous treatment effects. Section 4 discusses modified estimators that allow to recover of the ATE of observed and true switching units. Section 5 proposes a testing procedure to detect misspecification in the form of misclassification or anticipation as well as pre-trend violations. Section 6 examines the finite-sample behavior of the tests and the proposed estimators. Finally, Section 7 presents the empirical application, and Section 8 concludes. All proofs are contained in the Appendix.
2 Setup
[…]
The observation rule in our setting is given by:
$Y_{i,g,t}=Y_{i,g,t}(D_{i,g,t}^{*})=D_{i,g,t}^{*}Y_{i,g,t}(1)+\left(1-D_{i,g,t}^{*}\right)Y_{i,g,t}(0)$
We further define averages of observed treatment indicators at the group-period level:
$D_{g,t}$ $:=\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}D_{i,g,t},\ \mathbf{{D}}_{g}:=\left({D}_{g,t}\right)_{t\in\{1,\ldots,T\}},\ \mathbf{{D}}:=\left({D}_{g,t}\right)_{(g,t)\in\{1,\ldots,G\}\times\{1,\ldots,T\}},$
with analogous definitions other quantities such as $D_{g,t}^{*}$, $Y_{g,t}$ and $Y_{g,t}(0)$. Since $D_{i,g,t}\neq D_{i,g,t}^{*}$ may occur with positive probability, the above setup explicitly accommodates scenarios in which individuals experience the treated potential outcome $Y_{i,g,t}(1)$ despite having an observed treatment status $D_{i,g,t}=0$, or vice versa. We refer to these discrepancies as instances of misclassification. [^2] As noted earlier, misclassification can arise when exact treatment timing is unobserved or when treatment status is indirectly inferred. Furthermore, in Section 3, we demonstrate that this notation readily captures anticipatory effects, where units experience $Y_{i,g,t}(1)$ prior to the observed treatment.
We make the following assumptions:
Assumption 1 (Balanced Panel of Groups).
For each $(g,t)\in\{1,\ldots,G\}\times\{1,\ldots,T\}:$
$N_{g,t}>0.$
Assumption 2 (Sharp Design).
For each $(g,t)\in\{1,\ldots,G\}\times\{1,\ldots,T\}$ and $i\in\left\{1,\ldots,N_{g,t}\right\}:$
$D_{i,g,t}={D}_{g,t}.$
Assumption 3 (Mean Independence).
For each $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$:
$\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}\right]=\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}\right].$
Assumption 4 (Strong Exogeneity).
For each $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$:
$\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{{D}}_{g}\right]=\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right].$
Assumption 5 (Parallel Trends).
For each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
$\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]=\mathrm{E}\left[{Y}_{k,t}(0)-{Y}_{k,t-1}(0)\right].$
Similar to De Chaisemartin and D’Haultefoeuille (2020), we replace the independent groups assumption with mean independence, since the presence of stable groups (as defined below) inherently violates independence across groups (specifically, observed treatment status cannot be independent across groups in this case). Moreover, our sharp design assumption (Assumption 2) is stated with respect to the observed treatment status $D_{i,g,t}$, thus explicitly accommodating scenarios in which researchers mistakenly assume a sharp design based solely on observed data. Importantly, we allow for violations of the sharp design at the level of the latent treatment status $D_{i,g,t}^{*}$, recognizing that misclassification may occur within specific groups and periods. Furthermore, mean independence and strong exogeneity are defined with respect to the observed treatment status $\mathbf{D}_{g}$. Finally, similar to De Chaisemartin and D’Haultefoeuille (2020), we maintain a common trends assumption as our core identification assumption throughout the paper. [^3] See De Chaisemartin and D’Haultefoeuille (2020) for a discussion of the plausibility of the common trends assumption in settings with heterogeneous treatment effects.
We provide sufficient conditions for our assumptions to hold:
Proposition 1 (Characterizing Mean Independence and Strong Exogeneity with Observed Treatment Status).
Suppose that mean independence and strong exogeneity hold with respect to the actual treatment status $\mathbf{D}^{*}$:
$\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}^{*}\right]=\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right],$ (1)
for all $(g,t)\in\{1,\ldots,G\}\times\{1,\ldots,T\}$, and
$\mathrm{E}\left[Y_{g,t}(0)-Y_{g,t-1}(0)\mid\mathbf{D}_{g}^{*}\right]=\mathrm{E}\left[Y_{g,t}(0)-Y_{g,t-1}(0)\right],$ (2)
for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$. Then Assumptions 3 and 4, stated with respect to the observed treatment status $\mathbf{D}$, hold if the following are satisfied for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$:
  1. 1.
    Nondifferential misclassification:
    $\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*},\mathbf{D}_{g}\right]$ $=\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*}\right],$ (3)
    $\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}^{*},\mathbf{D}\right]$ $=\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}^{*}\right].$ (4)
  2. 2.
    Conditionally independence:
    $\mathbf{D}_{g}^{*}\perp\mathbf{D}_{-g}\mid\mathbf{D}_{g},$ (5)
    where $\mathbf{D}_{-g}:=\left(D_{k,t}\right)_{(k,t)\in\{1,\ldots,g-1,g+1,\ldots,G\}\times\{1,\ldots,T\}}$.
The proposition bears some resemblance to Lemma 1 of Denteh and Kedagni (2022). It shows that if misclassification is “nondifferential”, meaning it is independent of potential outcomes given the true treatment status (e.g., Lewbel, 2007a; Denteh and Kedagni, 2022), and if misclassification within each group is conditionally independent of the observed treatment statuses of other groups, then mean independence and strong exogeneity with respect to the true treatment status $\mathbf{D}_{g}^{*}$ imply the analogous assumptions with respect to the observed treatment status $\mathbf{D}_{g}$.
To further illustrate the mean independence assumption, note that (3) is equivalent to
$\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*}=d,\mathbf{D}_{g}=d^{\prime}\right]=\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*}=d,\mathbf{D}_{g}=d^{\prime\prime}\right],$
for all $d\in[0,1]$ and $d^{\prime},d^{\prime\prime}\in\{0,1\}^{T}$, provided the expectations are well-defined. This condition states that, given the actual treatment status, potential nontreatment outcomes exhibit the same trends regardless of whether the treatment status is correctly classified or misclassified. We also note that (5) is weaker than the joint independence $(\mathbf{D}_{g}^{*},\mathbf{D}_{g})\perp\mathbf{D}_{-g}$.
We also assume staggered adoption designs, where observed and true treatment represents “absorbing states”:
Assumption 6 (Staggered Adoption Designs).
For each $g\in\{1,\cdots,G\}$ and $t\in\{2,\ldots,T\}$:
$D_{g,t}\geq D_{g,t-1}.$
Assumption 6 is imposed solely for simplifying the notation. For designs involving more general switching patterns, similar assumptions about misclassification can be made for different types of switchers.
3 Bias Analysis
3.1 Homogenous Treatment Effects
In the absence of misclassification, it is well known that the standard TWFE estimator is generally biased in staggered adoption designs when treatment effects are heterogeneous. A notable exception is the case of homogeneous treatment effects, where
$Y_{i,g,t}(1)-Y_{i,g,t}(0)=:\Delta$ (6)
[…]
$\mathrm{E}\left[\widehat{\beta}_{fe}\right]=\mathrm{E}\left[\frac{\sum_{(g,t)}\frac{N_{g,t}}{N_{1}}\varepsilon_{g,t}D_{g,t}\mathrm{E}\left[\Delta\left|\mathbf{D}\right.\right]}{\sum_{(g,t):D_{g,t}=1}\frac{N_{g,t}}{N_{1}}D_{g,t}\varepsilon_{g,t}}\right]=\Delta,$
where $\varepsilon_{g,t}$ is the estimated residual from the unit-level regression of $D_{g,t}$ on group and period specific fixed effects. Unfortunately, this is no longer true in the presence of misclassification as the following result demonstrates:
Proposition 2 (Two-Way Fixed Effects).
Suppose Assumptions 1-5 and Equation (6) hold, as well as $\mathrm{E}\left[|\widehat{\beta}_{fe}|\right]<\infty$. Then, the TWFE estimand $\mathrm{E}\left[\widehat{\beta}_{fe}\right]$ is given by:
$\mathrm{E}\left[\widehat{\beta}_{fe}\right]=\Delta\mathrm{E}\left[\frac{\sum_{(g,t)}\frac{N_{g,t}}{N_{1}}\varepsilon_{g,t}\left(\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{Pr}\left(D_{i,g,t}^{\ast}=1\mid\mathbf{D}\right)\right)}{\sum_{(g,t)}\frac{N_{g,t}}{N_{1}}\varepsilon_{g,t}D_{g,t}}\right]$
The above result illustrates that even under the restrictive assumption of homogeneous treatment effects, misclassification generally introduces bias into the TWFE estimator. Whether this bias results in attenuation or amplification depends on the behavior of the misclassification probability $\mathrm{Pr}(D_{i,g,t}^{*}=1\mid\mathbf{D})$.
In the subsequent subsections, we extend this analysis to show that misclassification induces bias not only in the TWFE estimand under homogeneous effects, but also in more general settings with heterogeneous treatment effects, even when considering estimands specifically designed for the latter scenario.
3.2 ATE of Observed Switchers
To formalize the source of bias in what follows, we focus on a specific form of misclassification that occurs in the observed untreated period(s) before switching to treatment. This type of measurement error is common in applied settings where the precise timing of treatment is not observable due to data limitations. For instance, in a study of the causal effects of capital punishment reform on jury behavior, Bindler and Hjalmarsson (2018) define the “treatment year” as the first year with zero recorded death sentences, despite acknowledging that the actual legal change may have taken place earlier. In a staggered adoption setting, such coding can lead to systematic misclassification of pre-treatment periods in treated groups. The following assumption captures this structure of misclassification around the time of observed treatment adoption:
Assumption 7 (Patterns of Misclassification).
For each $(i,g,t)\in\{1,\cdots,N_{g,t}\}\times\{1,\ldots,G\}\times\{1,\ldots,T\}$ and $(j,g^{\prime},t^{\prime})\in\{1,\cdots,N_{g,t-1}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
  1. 1.
    $\mathrm{Pr}\left(D_{i,g^{\prime},t}^{*}=1\mid D_{g^{\prime},t}=1\right)=1$.
  2. 2.
    $\mathrm{Pr}\left(D_{j,g^{\prime},t^{\prime}-1}^{*}=0\mid D_{g^{\prime},t^{\prime}}=0,D_{g^{\prime},t^{\prime}-1}=0\right)=1$.
Assumption 7 requires that the unit-level observed treatment status is correctly classified in two situations: (i) periods in which a group is observed to receive treatment, and (ii) observed untreated periods that are followed by another observed untreated period. This assumption restricts misclassification to periods immediately preceding observed treatment, a simplification made for tractability and could be relaxed in more general settings. [^5] Note however that this formulation allows for misclassification to occur in the comparison group during the “treatment year”, which is particularly relevant when using “not-yet-treated” groups as controls. For instance, suppose a group is observed to be untreated in periods $t-1$ and $t$ (${D}_{g,t-1}=0$, ${D}_{g,t}=0$), but receives treatment in period $t+1$ (${D}_{g,t+1}=1$). Then, misclassification may arise in period $t$. In the context of Bindler and Hjalmarsson (2018), this means that units from the year before the “treatment year”, defined as the first year with zero death sentences, may be misclassified due to uncertainty about the timing of the legal reform, while units from other years are correctly classified.
Let $N_{S}:=\sum_{(g,t):t\geq 2,D_{g,t}>D_{g,t-1}}N_{g,t}$ denote the number of observations in the observed switching-to-treatment cells. In staggered adoption designs, this is the only type of observed switchers. To keep the analysis tractable, we focus on this staggered case and do not consider more general switching patterns, although such extensions are conceptually straightforward at the cost of more complex assumptions and arguments. We begin by defining a first causal parameter of interest:
$\delta^{S}$ $:=\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(i,g,t):t\geq 2,D_{g,t}\neq D_{g,t-1}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$=\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(i,g,t):t\geq 2,D_{g,t}>D_{g,t-1}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right],$
where the second equality uses Assumption 6. This quantity represents the ATE for units in the observed switching-to-treatment cells, where a group is observed to initiate or to receive treatment, across all switching periods among eventually treated groups.
We consider $\delta^{S}$ a natural starting point, as it coincides with the ATE of switchers in De Chaisemartin and D’Haultefoeuille (2020) when there is no misclassification. [^6] Note that, in the absence of misclassification, the ATE of switchers can in fact be viewed as a weighted average (weighted by the occurence probability of each group) of $\mathrm{ATE}(t,t)$s across time, where $\mathrm{ATE}(t,t)$ is the ATE in period $t$ of groups experiencing treatment at time $t$ (cf. Callaway and Sant’Anna, 2021). Even under misclassification, it remains a meaningful quantity, since it refers to a clearly defined subpopulation for which treatment is observed to begin in the data and serves as a useful benchmark. However, due to the patterns described in Assumption 7, some observed switching cells align with the actual switch, while others correspond to periods after treatment has already started. As a result, $\delta^{S}$ averages treatment effects from a mixture of true switchers and units who actually switched to treatment earlier than observed. While we demostrate the estimator proposed by De Chaisemartin and D’Haultefoeuille (2020) is generally biased in this case, we also show how to adjust it in the presence of misclassification. In Subsection 3.3, we also introduce the ATE of true switchers, an arguably more meaningful target of interest, which we aim to identify and estimate in Section 4.
Before introducing the estimator for $\delta^{S}$, we impose an additional assumption on the existence of comparison groups, following De Chaisemartin and D’Haultefoeuille (2020):
Assumption 8 (Existence of Comparison Groups).
For each $t\in\{2,\ldots,T\}$ and $g\in\{1,\ldots,G\}$ such that
$D_{g,t-1}=0,\ D_{g,t}=1,$
there exists $g^{\prime}\in\{1,\ldots,G\}$ with $g^{\prime}\neq g$ such that
$D_{g^{\prime},t-1}=D_{g^{\prime},t}=0.$
Let $N_{1,0,t}:=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}$ and $N_{0,0,t}:=\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}$ denote the number of observations in, respectively, the observed switching-to-treatment cells and the comparison cells at time $t$. Define the estimator
$\widehat{\operatorname{DID}}:=\frac{1}{N_{S}}\sum_{t=2}^{T}N_{1,0,t}\widehat{\operatorname{DID}}_{t},$ (7)
where
$\widehat{\operatorname{DID}}_{t}:=\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right),$
is set to zero when there is no switching-to-treatment group ($D_{g,t}=1$, $D_{g,t-1}=0$), or no comparison group ($D_{g,t}=0,D_{g,t-1}=0$). Intuitively, $\widehat{\operatorname{DID}}$ first computes a time-specific difference-in-differences $\widehat{\operatorname{DID}}_{t}$ by comparing outcome changes between observed switching and comparison groups, and then aggregates all these two-by-two comparisons across all time periods. In staggered adoption designs, this estimator coincides with the estimator labelled as $\widehat{\mathrm{DID}}_{M}$ suggested by De Chaisemartin and D’Haultefoeuille (2020).
Heuristically, the DID estimator based on the observed treatment status may fail to identify the ATE of units in the observed switching-to-treatment cells. This is because, in staggered adoption designs, the identifying parallel trends assumption concerns only the evolution of the potential nontreatment outcome $Y_{i,g,t}(0)$. However, under misclassification, the observed just-before-switching cells are contaminated by a mix of potential nontreatment and treatment outcomes. A similar issue arises in the comparison cells.
The following proposition formalizes this argument and shows that the DID estimand, $\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ can be decomposed into three components: (1) the ATE of units in the observed switching-to-treatment cells, (2) a bias term that relates to the ATE of the misclassified, actually treated units in the observed just-before-switching cells, and (3) a second bias term that relates to the ATE of the misclassified, actually treated units in the comparison cells.
Proposition 3 (Characterizing DID estimand).
Suppose Assumptions 1-8 hold. Then, the DID estimand $\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ can be decomposed as:
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ $=\delta^{S}-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}D_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\qquad\ -\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{1,0,t}}{N_{0,0,t}}D_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$ (8)
The first term in (8) is the ATE of units in the observed switching-to-treatment cells. The second term captures bias arising from actually treated units in the observed just-before-switching cells, who are misclassified as untreated. It disappears if such misclassification is absent, i.e., ${D}_{i,g,t-1}^{*}=0$ for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$ such that $D_{i,g,t-1}=0$ and $D_{i,g,t}=1$. To facilitate interpretation, assume $\frac{N_{g,t}}{N_{g,t-1}}=1$ for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$, a requirement that is satisfied when the data are a balanced panel. then this bias can be interpreted as the reweighted ATE of the misclassified, actually treated units in the observed just-before-switching cells,
$\mathrm{E}\left[w\left(\mathbf{D}\right)\frac{1}{N_{1,1;1,0}}\sum_{(i,g,t):t\geq 2,D_{i,g,t}^{*}=1,D_{i,g,t-1}^{*}=1,D_{i,g,t}=1,D_{i,g,t-1}=0}Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right],$
where $N_{1,1;1,0}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{D_{i,g,t}^{*}=1,D_{i,g,t-1}^{*}=1,D_{i,g,t}=1,D_{i,g,t-1}=0\right\}$, and $w\left(\mathbf{D}\right):=\frac{N_{1,1;1,0}}{N_{S}}$ is the weighting factor and between zero and one with probability one. Similarly, the third is the bias due to the misclassification of the actually treated units in the comparison cells of the observed switching-to-treatment cells, which disappears if ${D}_{i,g,t}^{*}=0$ for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$ such that $D_{i,g,t-1}=0$ and $D_{i,g,t}=0$. Loosely speaking, it relates to the ATE of the misclassified, actually treated units in comparison cells of the observed switching-to-treatment cells, however, we do not interpret it further due to the complexity of the weighting. [^7] If misclassification is not only present in periods just before observed switching-to-treatment cells, one can show that a term analogous to the third term in (8) for comparison groups appears also for period $t-1$ in the bias expression.
3.2.1 Illustration Example
For illustration, consider a stylized scenario with a balanced panel satisfying $N_{g,t}=N_{g,t-1}$ for all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$ and treatment effects are homogeneous across groups and over time:
${Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)=\Delta\in\mathbb{R},$
for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$. Suppose misclassification is homogeneous within all observed just-before-switching cells:
$\mathrm{E}\left[{D}_{i,g,t-1}^{*}\mid D_{i,g,t}=1,D_{i,g,t-1}=0,\mathbf{D}_{-(i,g,t,t-1)}\right]=p\in\mathbb{R},$
for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$, where
$\mathbf{D}_{-(i,g,t,t-1)}:=\left({D}_{i,g,t}\right)_{(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{1,\ldots,T\}\setminus\{(i,g,t-1),(i,g,t)\}}$
[…]
$\mathrm{E}\left[{D}_{i,g,t}^{*}\mid D_{i,g,t}=0,D_{i,g,t-1}=0,\mathbf{D}_{-(i,g,t,t-1)}\right]=0,$
for all $(i,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$. Then, the second term of the decomposition formula (8) becomes
$\quad\ \mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t}}D_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$=\Delta\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\mid\mathbf{D}\right]\right]$
$=\Delta\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\mid D_{i,g,t}=1,D_{i,g,t-1}=0,\mathbf{D}_{-(i,g,t,t-1)}\right]\right]$
$=\Delta p,$
while the last term disappears, so the bias boils down to attenuation bias:
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]=\Delta(1-p).$
3.3 ATE of True Switchers
As an alternative quantity of interest, we also discuss the ATE of true or actual switching units in the period of the recorded switch or prior. We refer to this quantity as ATE of True Switchers. To this end, recall the unobserved switching variable $S^{\ast}_{i,g,t}$ is one if a person with $D_{i,g,t}^{\ast}=1$ actually switched in that period $t$, and zero otherwise.
Thus, let $N_{S^{*}}:=\sum_{(i,g,t):t\geq 2}\mathbb{I}\left\{S^{*}_{i,g,t}=1\right\}$ denote the number of true switching units. The ATE of true switching units is given by:
$\delta^{S^{*}}$ $:=\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(i,g,t):t\geq 2,S_{i,g,t}^{*}=1}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right].$
As outlined in the introduction, $\delta^{S^{\ast}}$ measures the ATE of all switching-to-treatment units at the time when they actually switch status, across all switching periods among eventually treated groups in the data. This parameter may for instance be of interest when the initial impact of treatment is of particular policy relevance.
We obtain the following result for the identification of the ATE of True Switchers:
Proposition 4 (DID for True Switchers).
Suppose Assumptions 1-8 hold. Assume that for every $t\geq 2$ and $g\in\{1,\ldots,G\}$ such that $D_{g,t}>D_{g,t-1}$, it holds for all $(i,j)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,N_{g,t-1}\}$ that:
$\mathrm{Pr}\left({S}_{j,g,t-1}^{\ast}=1\mid D_{g,t}=1,D_{g,t-1}=0\right)+\mathrm{Pr}\left({S}_{i,g,t}^{\ast}=1\mid D_{g,t}=1,D_{g,t-1}=0\right)=1.$
On the contrary, for all $t\geq 2$ and $g\in\{1,\ldots,G\}$ such that $D_{g,t}=D_{g,t-1}$, it holds for all $(i,j)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,N_{g,t-1}\}$ that:
$\mathrm{Pr}\left({S}_{j,g,t-1}^{\ast}=1\mid D_{g,t}=D_{g,t-1}\right)+\mathrm{Pr}\left({S}_{i,g,t}^{\ast}=1\mid D_{g,t}=D_{g,t-1}\right)=0.$
Then, the DID estimand $\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ can be decomposed as:
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ $=\delta^{S^{\ast}}$
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$ (9)
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\left(\frac{N_{1,0,t}}{N_{0,0,t}}-1\right){D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$
Proposition 4 provides a decomposition of the bias of the DID estimand. The first two terms in (9) after $\delta^{S^{\ast}}$ are due to the fact that in the observed switching periods some individuals may in actuality already have undergone treatment in the earlier period and are thus not true switchers. In fact, the third, fourth, and fifth term also arise as there are already switchers in the “pre-treatment” periods prior to the recorded treatment periods. Finally, the last three terms appear because individuals from the comparison groups may actually also be switchers if their treatment occurs in a period just after the comparison.
[…]
Remark 1: Similar to the case of anticipation, our results and methodology are also applicable in setups with spillover effects. For instance, if a policy is enacted at the regional or state level, then units residing in a neighboring geographical location might experience $Y_{i,g,t}(1)$ even though the observed treatment status is recorded as $D_{i,g,t}=0$, and bias results similar to the ones in Propositions 3 and 4 can be derived. As an illustrative example, consider the legalization of recreational marijuana consumption for adults in the U.S., where different states adopted a legalization for recreational usage in different years (e.g., the state of New York in 2021). To the extent that (some) residents from neighboring states without a legalization such as Pennsylvania could “benefit” from the legal purchasing possibilities in New York by crossing the state border, the DID estimator $\widehat{\operatorname{DID}}$ from (7) will exhibit a bias with respect to the ATE of Observed and True Switchers, respectively.
4 Estimation of ATEs
In this section, we provide modified versions of the $\widehat{\mathrm{DID}}$ estimator to recover both the ATE of Observed Switchers, $\delta^{S}$, and the ATE of True Switchers, $\delta^{S^{\ast}}$. We begin with the former.
4.1 ATE of Observed Switchers
Let $N_{1,0,t}^{[1]}:=\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}$ and $N_{0,0,t}^{[1]}:=\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}$ and define
$\widehat{\mathrm{DID}}^{S}=\frac{1}{N_{S}}\sum_{t\geq 2}\widehat{\mathrm{DID}}^{S}_{t},$ (10)
where
$\widehat{\mathrm{DID}}^{S}_{t}=N_{1,0,t}\widehat{\operatorname{DID}}_{t}+N_{1,0,t}\widehat{\operatorname{DID}}_{t}^{*}+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\widehat{\operatorname{DID}}_{t}^{**}$
with
$\widehat{\operatorname{DID}}_{t}^{*}$ $=\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-2}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-2}\right)$
$\widehat{\operatorname{DID}}_{t}^{**}$ $=\frac{1}{N_{1,0,t}^{[1]}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right)-\frac{1}{N_{0,0,t}^{[1]}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right).$
Similar to the definition of $\widehat{\operatorname{DID}}_{t}$, each of the terms $\widehat{\operatorname{DID}}_{t}^{*}$ and $\widehat{\operatorname{DID}}_{t}^{**}$ is set to zero whenever the corresponding treatment or comparison group does not exist. In particular, we set $\widehat{\operatorname{DID}}_{2}^{*}=0$ and $\widehat{\operatorname{DID}}_{T}^{**}=0$, as no observations are available beyond periods 111 and $T$, respectively.
Heuristically, the estimator $\widehat{\mathrm{DID}}^{S}$ in (10) corrects the original $\widehat{\mathrm{DID}}$ estimator for the two bias terms identified in Proposition 3, using appropriate sample counterparts. Specifically, $\widehat{\operatorname{DID}}_{t}^{*}$ is a “backward-looking correction”: it compares outcome changes from $t-2$ to $t-1$, to adjust for bias due to misclassified treated units in the treatment group at $t-1$. In contrast, $\widehat{\operatorname{DID}}_{t}^{**}$ is a “forward-looking correction”: it compares outcome changes from $t-1$ to $t$ between units still observed as untreated at $t$, but differing in whether they begin treatment at $t+1$, to adjust for bias due to misclassified treated units in the comparison group at $t$. Together, these corrections yield an unbiased estimator of the ATE of observed switchers, $\delta^{S}$.
Since our correction relies on outcome information from both periods preceding and following the observed switch, we impose the following additional conditions:
Assumption 9 (No Misclassification in Never Treated).
For each $(i,g)\in\{1,\cdots,N_{g,T}\}\times\{1,\cdots,G\}$:
$\mathrm{Pr}\left(D_{i,g,T}^{*}=0\mid D_{g,T}=0\right)=1.$
Assumption 10 (No Observed Switch in Initial Period).
For each $g\in\{1,\cdots,G\}$:
$\Pr\left(D_{g,1}=D_{g,2}\right)=1.$
With staggered adoption designs and the misclassification patterns described in Assumption 7, Assumption 9 ensures that the never treated groups are always correctly classified. If a group indeed begins to receive treatment at $T+1$, it could be misclassified at time $T$, so that this condition is violated. In such cases, the researcher may address the issue by collecting data for one additional period, omitting the final period from analysis, or excluding the affected group (particularly, if the group is not central to the research question and sufficient comparison groups remain). On the other hand, Assumption 10 places a restriction on the data to rule out observed switches in the initial period. Combined with Assumption 7, it ensures that no misclassification occurs at $t=1$, which would otherwise be uncorrectable due to the absence of pre-period information.
We now establish that $\widehat{\mathrm{DID}}^{S}$ provides an unbiased estimator of $\delta^{S}$:
Proposition 5 (DID for Observed Switchers).
Suppose Assumptions 1-10 hold. Then:
$\mathrm{E}\left[\widehat{\mathrm{DID}}^{S}\right]$ $=\delta^{S}.$
4.2 ATE of True Switchers
Since $\delta^{S^{\ast}}$, the ATE of true switchers, constitutes arguably a more interesting quantity of interest than $\delta^{S}$, we next propose an estimator to recover the second target parameter of interest. To this end, recall the definition of $\delta^{S^{\ast}}$ as:
$\delta^{S^{\ast}}=\mathrm{E}\left[\frac{1}{N_{S^{\ast}}}\sum_{(i,g,t):t\geq 2,S^{\ast}_{i,g,t}=1}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$
We construct an estimator of $\delta^{S^{\ast}}$ as:
$\widehat{\mathrm{DID}}^{S^{\ast}}$ $=\frac{1}{\widehat{N}_{S^{\ast}}}\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}\widehat{\mathrm{DID}}^{S^{\dagger}}_{t-1}+\left(1-\widehat{\lambda}_{t-1}\right)\widehat{\mathrm{DID}}^{S}_{t}\right),$ (11)
where
$\widehat{\mathrm{DID}}^{S^{\dagger}}_{t-1}=N_{1,0,t-1}^{[1]}\widehat{\operatorname{DID}}_{t-1}^{**}+N_{1,0,t-1}^{[1]}\widehat{\operatorname{DID}}_{t-1}^{***}\quad\text{and}\quad\widehat{\lambda}_{t-1}\equiv\frac{\widehat{\operatorname{DID}}_{t-1}^{**}}{\widehat{\operatorname{DID}}_{t-1}^{**}+\widehat{\mathrm{DID}}_{t-1}^{***}},\quad$
and
$\widehat{N}_{S^{\ast}}\equiv\sum_{t\geq 2}\sum_{D_{g,t}=1,D_{g,t-1}=0}\left(\widehat{\lambda}_{t-1}N_{g,t-1}+(1-\widehat{\lambda}_{t-1})N_{g,t}\right),$
with the component:
$\widehat{\mathrm{DID}}_{t-1}^{***}=$ $\frac{1}{N_{1,0,t-1}^{[1]}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t-1}\left(Y_{g,t}-Y_{g,t-1}\right)-\frac{1}{N_{1,1,t-1}^{[1]}}\sum_{g:D_{g,t}=1,D_{g,t-1}=1}N_{g,t-1}\left(Y_{g,t}-Y_{g,t-1}\right),$
while $\widehat{\mathrm{DID}}^{S}_{t}$ and $\widehat{\mathrm{DID}}_{t}^{**}$ are given in (10). As before, we adopt the convention that $\widehat{\mathrm{DID}}_{t}^{***}$ is set to zero whenever any of the corresponding groups used therein does not exist in the data. We also set ${\operatorname{DID}}_{1}^{**}=0$, and set $\widehat{\lambda}_{t-1}=0$ when both $\widehat{\operatorname{DID}}_{t-1}^{**}$ and $\widehat{\operatorname{DID}}_{t-1}^{***}$ are zero. Heuristically, the estimator $\widehat{\mathrm{DID}}^{S^{\ast}}$ in (11) aggregates treatment effect estimates from both the observed just-before-switching and switching-to-treatment periods, weighted by the estimated probability of a true switch occurring in each. Specifically, under the restrictions specified below, $\widehat{\mathrm{DID}}^{S^{\dagger}}_{t}$ estimates the unscaled ATE in period $t-1$ for units who actually switched at time $t-1$ but are only observed to switch at time $t$, while $\widehat{\mathrm{DID}}^{S}_{t}$ estimates the unscaled ATE in period $t$ for units whose switch is both actual and observed at time $t$. Moreover, the weight $\widehat{\lambda}_{t-1}$ estimates the share of the former group, and $1-\widehat{\lambda}_{t-1}$ estimates that of the latter.
We require the following assumptions:
Assumption 11 (Patterns of Switch).
For each $(i,j,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,N_{g,t-1}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
  1. 1.
    $\mathrm{Pr}\left(S^{\ast}_{i,g,t}=0\mid D_{g,t}=1,\ D_{g,t-1}=1\right)=1$.
  2. 2.
    $\mathrm{Pr}\left(S^{\ast}_{j,g,t-1}=0\mid D_{g,t}=0,\ D_{g,t-1}=0\right)=1.$
In addition, for each $(i,g)\in\{1,\ldots,N_{g,T}\}\times\{1,\cdots,G\}$:
$\mathrm{Pr}\left(S_{i,g,T}^{*}=0\mid D_{g,T}=0\right)=1.$
Assumption 12 (Switcher Homogeneity).
For each $(i,j,g,t)\in\{1,\ldots,N_{g,t}\}\times\{1,\ldots,N_{g,t-1}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
  1. 1.
    $\mathrm{Pr}\left({D}_{j,g,t-1}^{\ast}=1\mid D_{g,t}=1,\ D_{g,t-1}=0\right)=\mathrm{Pr}\left({S}_{j,g,t-1}^{\ast}=1\mid D_{g,t}=1,\ D_{g,t-1}=0\right)=\lambda_{t-1}$.
  2. 2.
    $\mathrm{Pr}\left(S^{\ast}_{i,g,t}=1\mid D_{g,t}=1,\ D_{g,t-1}=0\right)+\mathrm{Pr}\left(S^{\ast}_{j,g,t-1}=1\mid D_{g,t}=1,\ D_{g,t-1}=0\right)=1.$
Assumption 13 (Treatment Effects Homogeneity).
For each $(i,g,t)\in\{1,\cdots,N_{g,t}\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$, it holds that:
$\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\mid D_{i,g,t}^{\ast},S_{i,g,t}^{\ast},\mathbf{D}_{g}\right]=\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\right].$
Assumption 11 characterizes the behavior of the switching variable $S_{i,g,t}^{\ast}$. It stipulates that $S_{i,g,t}^{\ast}$ depends on the immediate information contained in $D_{g,t}$ and $D_{g,t-1}$, not the entire observed treatment path of group $g$. Moreover, in staggered adoption designs, Assumption 11 complements Assumption 7 by imposing coherent restrictions in terms of the unobserved actual switching indicator $S^{\ast}_{i,g,t}$. Specifically, it rules out the possibility that a true switch, i.e., $S^{*}_{i,g,t}=1$, occurs within a run of treated or untreated periods, except at the first treated period or the last untreated period, respectively. The last part of the assumption parallels Assumption 9 and further rules out such a switch in the last untreated period among the never treated.
The first part of Assumption 12 imposes homogeneity on the switching process across groups and units. It conceptually links the latent treatment status $D^{*}_{i,g,t-1}$ with the switching indicator $S^{*}_{i,g,t-1}$, and requires that all units observed in groups that switch between periods $t-1$ and $t$ share the same probability $\lambda_{t-1}$ of having actually switched in the prior period. [^9] Recall that for notational simplicity, we focus on misclassification occurring only in the period immediately preceding the observed switch. However, the framework can be extended to accommodate more complex settings in which multiple pre-switch periods are affected by misclassification or anticipation. In such cases, one could allow for different pre-treatment misclassification probabilities, though at the cost of more cumbersome notation. The second part formalizes that this total switching probability is fully split between the switching-to-treatment and the just-before-switching periods.
Assumption 13 on the other hand essentially requires that expected unit-level treatment effects cannot be fully heterogeneous, in that they cannot depend on unit’s actual treatment status or switching status. Crucially, this assumption only restricts how such heterogeneity may correlate with actual treatment assignment and switching, while treatment effects may still differ across individuals, groups, or time.
To recover $\delta^{S^{\ast}}_{0}$ , we also require extensions of Assumptions 3-5 to potential outcomes under treatment:
Assumption 14 (Independent Groups).
The group-level vectors are mutually independent across $g$:
$\left(Y_{g,t}(0),Y_{g,t}(1),D_{g,t},(D_{i,g,t}^{\ast})_{i\in\{1,\cdots,N_{g,t}\}},\left(S_{i,g,t}^{\ast}\right)_{i\in\{1,\cdots,N_{g,t}\}}\right)_{1\leq t\leq T}$
Assumption 15 (Strong Exogeneity).
For each $(d,g,t)\in\{0,1\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
$\mathrm{E}\left[{Y}_{g,t}(d)-{Y}_{g,t-1}(d)\mid\mathbf{{D}}_{g}\right]=\mathrm{E}\left[{Y}_{g,t}(d)-{Y}_{g,t-1}(d)\right].$
Assumption 16 (Parallel Trends).
For each $(d,g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$:
$\mathrm{E}\left[{Y}_{g,t}(d)-{Y}_{g,t-1}(d)\right]=\mathrm{E}\left[{Y}_{k,t}(d)-{Y}_{k,t-1}(d)\right].$
Assumption 14 parallels Assumption 3 from De Chaisemartin and D’Haultefoeuille (2020) and imposes an independence assumption across groups (but not homogeneity). It replaces Assumption 3 and extends it (through the independence) to outcomes in the presence of treatment $Y_{i,g,t}(1)$. Similarly, Assumptions 15 and 16 also represent extensions of Assumptions 4 and 5 to potential outcomes under treatment. They are needed to identify the (unobserved) share of true switchers in a given switching-to-treatment cell, which in turn relies on a comparison with groups that have already switched to treatment in earlier periods. In fact, since standard staggered adoption settings without misspecification do not require such assumptions, these conditions may be viewed as the “price to pay” for not observing the actual switching status. We note however that more general DiD set-ups that also allow for “reverse switchers”, i.e. units switching status back from treatment to no treatment, also rely on Assumptions 14-16 for identification (cf. De Chaisemartin and D’Haultefoeuille, 2020; Fabre, 2023).
Finally, since the exact period of switching-to-treatment is unknown at the individual level and we may only identify the fraction of switchers in a given period in an asymptotic framework, we focus on an estimator that is consistent for $\delta^{S^{\ast}}$ in this section. To establish consistency, we also require a couple of technical conditions concerning limiting quantities:
Assumption 17 (Existence of Comparison Groups - 2).
For each $g\in\{1,\ldots,G\}$ and $t\in\{2,\ldots,T\}$ such that
$\Pr\left(D_{g,t}=1,D_{g,t-1}=0\right)>0,$
implies that $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ and $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{0,0,t}}{G}\right]>0$. Moreover, it also implies that $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{j,0,t}^{[1]}}{G}\right]>0$ and $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{j,0,t-1}^{[1]}}{G}\right]>0$ as well as $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,1,t-1}^{[1]}}{G}\right]>0$. Finally, for at least one $t\in\{2,\ldots,T\}$, it holds that $\Pr\left(D_{g,t}=1,D_{g,t-1}=0\right)>0$.
Assumption 18 (Existence of Moments).
$\sup_{(g,t)}N_{g,t}<+\infty$ and $\sup_{(d,i,g,t)}\mathrm{E}\left[Y_{i,g,t}(d)^{4}\right]<+\infty$. Moreover, let $\lim_{G\rightarrow\infty}\mathrm{E}[\mathbf{V}]$ and $\mathbf{\Sigma}_{\mathbf{V}}$ defined in (20) and (21) of Appendix 9.1 exist.
Assumption 17 is an asymptotic extension of Assumption 8, that does not only involve comparison groups of the form $D_{g,t}=0$, $D_{g,t-1}=0$, but also all other groups involved in the construction of $\widehat{\mathrm{DID}}^{S^{\ast}}$. Assumption 18 requires that the fourth moments of the potential outcomes $Y_{i,g,t}(d)$ are uniformly bounded across all indexes. It also assumes that the cell sizes $N_{g,t}$ are uniformly bounded. Finally, the assumption ensures the existence of the limiting expectation vector $\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ and the asymptotic covariance $\boldsymbol{\Sigma}_{\mathbf{V}}$, both of which are required for establishing the asymptotic distribution of the estimator.
Considering an asymptotic framework where the number of groups $G$ tends to infinity (with the number of time periods $T$ fixed), we have the following asymptotic result:
Proposition 6 (DID for True Switchers).
Suppose that Assumptions 1-2, 6-7 and 9-18 hold. Moreover, it holds that $\lim_{G\rightarrow\infty}\Pr\left(\left|\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{**}\right]+\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{***}\right]\right|>0\right)>0$ for all $t\in\{2,\ldots,T\}$. Then:
$\sqrt{G}\left(\widehat{\mathrm{DID}}^{S^{\ast}}-\delta^{S^{\ast}}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\sigma_{\delta^{S^{\ast}}}^{2})$
with $\sigma_{\delta^{S^{\ast}}_{0}}^{2}:=\mathbf{J}_{f}\left(\mathbf{V}^{\infty}\right)\boldsymbol{\Sigma}_{\mathbf{V}}\mathbf{J}_{f}\left(\mathbf{V}^{\infty}\right)^{\prime}$, where $\mathbf{V}$ is defined in (20) and $\mathbf{V}^{\infty}=\lim_{G\rightarrow\infty}\mathrm{E}\left[\mathbf{V}\right]$.
Proposition 6 establishes the limiting distribution of $\widehat{\mathrm{DID}}^{S^{\ast}}$. Note that the condition:
$\lim_{G\rightarrow\infty}\Pr\left(\left|\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{**}\right]+\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{***}\right]\right|>0\right)>0$ (12)
is a technical condition that ensures a well-defined limit. It requires the existence of a non-zero treatment effect prior to treatment, but can be weakened by introducing a suitable trimming function into the estimator $\widehat{\mathrm{DID}}^{S^{\ast}}$. Specifically, denote by $\mathcal{T}_{DID,G}$ the set of periods such that
$\left|\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{**}\right]+\mathrm{E}\left[\widehat{\operatorname{DID}}_{t-1}^{***}\right]\right|>c_{G}$
for some $c_{G}\rightarrow 0$ as $G\rightarrow\infty$. Then, the estimator in (11) can be modified as:
$\widehat{\mathrm{DID}}^{S^{\ast}}$ $=\frac{1}{\widehat{N}_{S^{\ast}}}\sum_{t\geq 2}\mathbb{I}\left\{t\in\mathcal{T}_{DID,G}\right\}\left(\widehat{\lambda}_{t-1}\widehat{\mathrm{DID}}^{S^{\dagger}}_{t-1}+\left(1-\widehat{\lambda}_{t-1}\right)\widehat{\mathrm{DID}}^{S}_{t}\right).$ (13)
Under suitable rate conditions, the modified estimator in (13) is a consistent estimator for the ATE of True Switchers for all periods in $\mathcal{T}_{DID,G}$. Alternatively, a continuous, differentiable function could also be used for trimming (e.g., Gutknecht, 2016). Finally, note that the testing procedure outlined in the next section allows to test for the presence of (non-zero) anticipation or misclassification effects, and may thus also inform the researcher whether (12) may be an appropriate assumption or not in a given setting.
5 Specification Tests
As illustrated in the previous sections, not accounting for misclassification or anticipation in the DID estimator from Equation (7) leads to a bias not only with respect to $\Delta$ in the case of homogeneous treatment effects, but also with respect to both $\delta_{S}$ and $\delta_{S^{\ast}}$, respectively. As a result, a modification of DID is required. However, in the absence of further information, it is not clear how to modify the aforementioned estimator when the researcher does not know how many periods prior to the observed switches are affected specifically. In addition, even if knowledge about affected periods was available, inspection of the proof of Proposition 3 suggests that it is generally difficult to disentangle the bias arising from misclassification or anticipation on the one hand and from a failure of parallel trends (PT) in Assumption 5 on the other. This is, however, important as the implications are different. In particular, while anticipation or misclassification in pre-switch periods can be addressed through a modified DID estimator using earlier periods for comparison (subject to data availability), a failure of strong exogeneity and thus PT is generally invalidating the DID approach for point identification and estimation altogether. [^10] Of course, it may still be possible to adopt a partial identification approach as in Rambachan and Roth (2023) using pre-treatment evidence to bound post-treatment outcome differences. The testing procedure may thus be more informative than standard, event-study type pre-trend tests that do not differentiate between these two different forms of misspecifications.
To address the aforementioned challenges, we propose a two-step testing procedure to detect misspecification due to anticipation effects or misclassifcation, and to guide researchers in the adaptation of the corresponding estimators. In particular, in a first step we aim to maximize confidence in the PT assumption by examining pre-treatment evidence for its validity comparing treated groups with not-yet switched comparison groups in time periods that are known not to be affected by misclassification. Conditional on passing this initial stage, we then conduct a second test for misclassification or anticipation by comparing the same groups using periods that could be misclassified. The outcome of this second test provides guidance on how to adapt the DID estimator to recover $\delta_{S}$ or $\delta_{S^{\ast}}$.
To implement this test procedure, recall that the parallel trends assumption naturally extends to other pre-treatment periods. That is, if for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{2,\ldots,T\}$, it holds that
$\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]=\mathrm{E}\left[{Y}_{k,t}(0)-{Y}_{k,t-1}(0)\right],$
then it also holds that:
$\mathrm{E}\left[{Y}_{g,t-1}(0)-{Y}_{g,t-l}(0)\right]=\mathrm{E}\left[{Y}_{k,t-1}(0)-{Y}_{k,t-l}(0)\right]$
for each $(g,k,t)\in\{1,\ldots,G\}\times\{1,\ldots,G\}\times\{(l+1),\ldots,T\}$ with $l\in\{2,\ldots,T-1\}$. Define the following quantities:
$\widehat{\mathrm{\tau}}_{1,0}(t,l)$
$=$ $\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right).$ (14)
This is a “pre-treatment” statistic comparing pre-treatment periods of switching groups with those same periods of not-yet switching groups. Also, let
$\widehat{\mathrm{\tau}}_{0,0}(t,l)$
$=$ $\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right).$ (15)
This is a “pre-treatment” statistic comparing pre-treatment periods of switching groups with those of not-yet switching groups. To understand the heuristic of the testing procedure, let
$\psi_{0,t-1,l}=\mathrm{E}\left[Y_{g,t-1}(0)-Y_{g,t-l}(0)\right].$
We note that in the presence of misspecification (i.e., anticipation or misclassification) and violation of PT, under Assumptions 1-4, 6-8, the expectation of $\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)$ (provided it exists) is equal to:
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}D_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$+\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\psi_{0,t-1,l}-\frac{1}{{N}_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\psi_{0,t-1,l}\right].$
On the other hand, for the second statistic, we only have a potential violation of the parallel trends part (by the assumption that misspecification only occurs in cells before the switch):
$\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}^{(l)}(t)\right]=$ $\mathrm{E}\left[\frac{1}{N_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\left(Y_{g,t-2}-Y_{g,t-l}\right)\right]$
$=$ $\mathrm{E}\left[\frac{1}{{N}_{1,0,t}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}-\frac{1}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\psi_{0,t-2,l}\right].$
This suggests a testing procedure for anticipation or misclassification effects based on moment equalities focusing on the probability limits in the sense of De Chaisemartin and D’Haultefoeuille (2020) or of Callaway and Sant’Anna (2021). Thus, denote:
$\mathrm{\tau}_{1,0}^{(l)}(t)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right]$
and accordingly:
$\mathrm{\tau}_{0,0}^{(l)}(t)\equiv\lim_{G\rightarrow\infty}\mathrm{E}\left[\widehat{\mathrm{\tau}}_{0,0}^{(l)}(t)\right]$
for every $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. The testing procedure is based on the following heuristic: if there is pre-treatment evidence suggesting a likely violation of PT, DID should generally not be used for point identification and estimation. On the other hand, if there is pre-treatment evidence supporting PT, we may actually proceed with DID and seek to design the estimator in such a way to avoid misspecification bias. Put differently, we aim to “maximize” our confidence in the PT assumption in a first testing step by looking at pre-treatment evidence in favor of PT, before turning to a misspecification test in a second step to obtain guidance on how to manipulate the estimator to avoid bias. In fact, building on pre-treatment evidence for PT is in line with more recent work by Rambachan and Roth (2023) who exploit pre-treatment PT violations as a way to infer about possible post-treatment violations. Moreover, the procedure is also motivated by the recent findings of De Chaisemartin and D’Haultefoeuille (2024) who show that conducting pre-treatment specification tests may, under the null of correct specification, generally only lead to conservative, but not to too liberal post-treatment inference. This in turn suggest that post-treatment inference carried out conditional on passing pre-treatment specification test, maintains asymptotic size control under the null of correct specification.
To formalize the testing procedure, we first define two distinct tests. The first is for the violation of PT. Specifically, we want to test:
$H_{0}^{PT}:\mathrm{\tau}_{0,0}^{(l)}(t)=0$
for all $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$, versus:
$H_{A}^{PT}:\mathrm{\tau}_{0,0}^{(l)}(t)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Similarly, we define the second null hypothesis to be tested as:
$H_{0}^{MC}:\mathrm{\tau}_{1,0}^{(l)}(t)=0$
for all $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$, versus:
$H_{A}^{MC}:\mathrm{\tau}_{1,0}^{(l)}(t)\neq 0$
for some $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. Note that the above tests can be carried out via moment equality tests. We will outline the test for $H_{0}^{MC}$ against $H_{A}^{MC}$ in what follows, but note that the same results hold for $H_{0}^{PT}$ against $H_{A}^{PT}$. In particular, we may use a statistic of the form:
$T_{1,0}^{s}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right)^{2},$ (16)
or
$T_{1,0}^{m}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t)\right|,$ (17)
where $\mathcal{C}_{S}$ is an index set $\mathcal{C}_{S}=\{(l,t):\;t\in\{(l+1),\ldots,T\};\;l\in\{2,\ldots,T-1\}\}$ with cardinality $\mathcal{C}_{S}$. For the second test, an analogous statistic can be used.
Remark 2: The tests above are designed to cover all pre-treatment periods provided there is only a finite number of them. However, in practice, as researchers may not want to use evidence from far away pre-treatment periods, we can restrict the number of pre-treatment periods used in the corresponding moment equality tests. At the same time, observe that the moment equality testing framework is more flexible than standard pre-trend tests that only test the significance of a (some) DiD parameter(s) relative to a single, fixed pre-treatment reference period. Specifically, the test statistics in (16) and (17) allow to include also comparisons of other combinations of pre-treatment periods. In practice, this may yield more power to detect specific violations of misclassification and PT in pre-treatment periods than what could be detected with only a fixed reference period.
Remark 3: If the researcher believes that also other pre-treatment periods are affected by misspecification, one may set a threshold, say $\kappa$, beyond which the researcher has certainty that no misclassification or anticipation is present so that the PT only test can be carried out.
We make the following assumptions:
Assumption 19 (Existence of Moments).
Let $\lim_{G\rightarrow\infty}\mathrm{E}[\mathbf{U}_{t,l}]$ and $\mathbf{\Sigma}_{\mathbf{U}}$ defined in (22) of Appendix 9.1 exist. Moreover, it holds that $\sup_{(d,g,t)}\mathrm{E}\left[Y_{g,t}(d)^{8+\delta}\right]<+\infty$ for some $\delta>0$.
Assumption 20 (Stable Groups).
For all $(g,t)\in\{1,\ldots,G\}\times\{2,\ldots,T\}$, we have that
$\Pr\left(D_{g,t}=1,D_{g,t-1}=0\right)>0$
implies $lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ and $lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{0,0,t}}{G}\right]>0$. Moreover, $lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ for at least one $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$.
Since the limiting distribution of our tests under the null hypothesis is non-standard, we used the nonparametric (block) bootstrap to derive critical values. To do so, let
$\mathbf{{Y}}_{g}:=\left({Y}_{g,t}\right)_{t\in\{1,\ldots,T\}}$
denote the time series vector of sample averages from each group $g$. Also, recall that:
$\mathbf{{D}}_{g}:=\left({D}_{g,t}\right)_{t\in\{1,\ldots,T\}}$
[…]
$T_{1,0}^{s,b}=\sum_{(t,l)\in\mathcal{C}_{S}}\left(\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)-\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t))\right)^{2},$
or
$T_{1,0}^{m,b}=\max_{(t,l)\in\mathcal{C}_{S}}\left|\sqrt{G}(\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)-\widehat{\mathrm{\tau}}_{1,0}^{(l)}(t))\right|.$
Here, $\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)$ denotes the bootstrap analog of the test statistic:
$\widehat{\mathrm{\tau}}_{1,0}^{(l),b}(t)=\frac{1}{N_{1,0,t}^{b}}\sum_{g:D^{b}_{g,t}=1,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-\frac{1}{N^{b}_{0,0,t}}\sum_{g:D^{b}_{g,t}=0,D^{b}_{g,t-1}=0}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right).$
with:
$N^{b}_{1,0,t}=\sum_{g\in G}\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}$
and
${N}^{b}_{0,0,t}=\sum_{g\in G}\mathbb{I}\{D^{b}_{g,t}=0,D^{b}_{g,t-1}=0\}N_{g,t}$
The critical values are then given by the $(1-\gamma)$, $\gamma\in(0,1)$, quantiles of the empirical bootstrap distributions of $T_{1,0}^{s,b}$ and $T_{1,0}^{m,b}$, respectively, over $B$ draws, say $c^{m}_{B,G}(1-\gamma)$ and $c^{s}_{B,G}(1-\gamma)$. In what follows, we focus on the test of $H_{0}^{MC}$, as the arguments for the test involving $T_{0,0}^{s}$ and $T_{0,0}^{m}$ are identical. Thus, define the $|\mathcal{C}_{S}|\times 1$ vector:
$\widehat{\boldsymbol{\tau}}_{1,0}=\left(\begin{array}[]{c}\widehat{\tau}_{1,0}(t,l)\\ \vdots\\ \widehat{\tau}_{1,0}(j,k)\end{array}\right)$ (18)
and similarly:
$\boldsymbol{\tau}_{1,0}=\left(\begin{array}[]{c}\tau_{1,0}(t,l)\\ \vdots\\ \tau_{1,0}(j,k)\end{array}\right)$ (19)
Finally, note that by Assumptions 19 and 20, we have with probability approaching 1 that there exists at least one $(t,l)$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]>0$ and so $\min\{N_{1,0,t},{N}_{0,0,t}\}>0$ for sufficiently large $G$. On the other hand, observe that for any element $(t,l)\in\mathcal{C}_{S}$ for which $\lim_{G\rightarrow\infty}\mathrm{E}\left[P_{1,0}(t,l)\right]=0$, the corresponding element does not contribute to the statistic $T_{1,0}^{s}$ for sufficiently large $G$, and does also not contribute to the asymptotic variance of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$. We therefore proceed by analyzing the case where all elements of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$ contribute to the asymptotic distribution, i.e. the case where $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{1,0,t}}{G}\right]>0$ and $\lim_{G\rightarrow\infty}\mathrm{E}\left[\frac{N_{0,0,t}}{G}\right]>0$ hold for all $(t,l)\in\mathcal{C}_{S}$ in Assumption 20. We have the following statement:
Proposition 7.
Assume that Assumptions 1-2, 4-7, and 14, 19, 20 hold for all $(t,l)\in\mathcal{C}_{S}$, and that $\boldsymbol{\Sigma}_{\mathbf{U}}$ from (22) is positive definite, and $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)$ defined in (27) has full column rank. Then:
(i) Under $H_{0}^{MC}$, it holds that:
$\lim_{G,B\rightarrow\infty}\Pr\left(T_{1,0}^{s}>c^{s}_{B,G}(1-\alpha)\right)=\alpha$
and
$\lim_{G,B\rightarrow\infty}\Pr\left(T_{1,0}^{m}>c^{m}_{B,G}(1-\alpha)\right)=\alpha.$
(ii) Under $H_{A}^{MC}$, it holds that:
$\lim_{G,B\rightarrow\infty}\Pr\left(T_{1,0}^{s}>c^{s}_{B,G}(1-\alpha)\right)=1$
and
$\lim_{G,B\rightarrow\infty}\Pr\left(T_{1,0}^{m}>c^{m}_{B,G}(1-\alpha)\right)=1.$
As remarked previously, we note that identical results for the test $H_{0}^{PT}$ against $H_{A}^{PT}$ as we also demonstrate in our Monte Carlo simulations.
Finally, even though both tests can be carried out idependently, we may also consider a decision rule for the joint execution of both tests that provides guidance for practicioners. In particular, there are three cases we would like to distinguish as they bear different implications for the empirical analysis (see above): a violation of PT in pre-treatment periods, possibly with misclassification ($H_{TV}$), PT in pre-treatment periods holding without misclassification ($H_{PTN}$), and PT in pre-treatment periods holding with misclassification ($H_{PTM}$). We have that:
$H_{TV}=H_{A}^{PT}$
$H_{PTN}=H_{0}^{PT}\cap H_{0}^{MC}$
$H_{PTM}=H_{0}^{PT}\cap H_{A}^{MC}$
The decision rule is as follows:
1. Test $H_{0}^{PT}$ vs. $H_{A}^{PT}$ at level $\alpha$. If $H_{0}^{PT}$ is rejected at level $\alpha$, stop and decide in favor of $H_{TV}$. If $H_{0}^{PT}$ is not rejected, proceed with step 2.
2. Test $H_{0}^{MC}$ vs. $H_{A}^{MC}$ at level $\gamma$. If $H_{0}^{MC}$ is not rejected at level $\gamma$, decide in favor of $H_{PTN}$ and use standard estimator. If $H_{0}^{MC}$ is rejected, decide in favor of $H_{PTM}$ and modify estimator.
6 Simulations
In order to assess the finite sample performance of the estimators and tests introduced in the previous sections, we conduct a Monte Carlo study. We start by comparing the two proposed estimators $\widehat{\mathrm{DID}}^{S}$ and $\widehat{\mathrm{DID}}^{S^{*}}$ to an estimator that does not account for misspecification in the treatment indicator. In a second step, we then analyze the rejection rates of the specification tests under various Data Generating Processes (DGP).
[…]
where $\Delta_{t}$ is the treatment effect, while $e_{g,t}$ denotes the idiosyncratic error term. We simulate both time constant and time-varying treatment effects. Specifically, the constant treatment effects are drawn from a normal distribution with a mean of 4 and a variance of 1, i.e. $\Delta\sim\mathcal{N}(4,1)$. The time varying treatment effects on the other hand are generated by
$\Delta_{t}=\Delta\,\left(0.2+(1.8-0.2)\,\frac{(t-1)}{(T-1)}\right),$
so that the mean treatment effect increases from 0.8 to 7.2 over time.
While the DGP is based on the true treatment status $D_{g,t}^{*}$, misclassification is induced via $D_{g,t}$, the recorded group-period treatment status. In line with the simplifying assumption from Sections 4 and 5, $D_{g,t}$ equals $D_{g,t}^{*}$ for all time periods but the period before treatment, in which 50% of the groups are assigned $D_{g,t}=0$, while $D_{g,t}^{*}=1$. Groups that enter treatment in the last time period are not misclassified.
Table 1 and Table 2 display the simulation results for the estimators presented in Section 4 in comparison to the estimator of De Chaisemartin and D’Haultefoeuille (2020) that does not account for misspecification in the treatment indicator. Table 1 contains the mean bias with respect to $\Delta$ or $\bar{\Delta}\equiv 1/14\sum_{t=2}^{15}\Delta_{t}$, respectively. [^12] Note that in population the benchmark ATE coincides with the ATE of true switchers since the misclassification or switching probabilities are constant over time in this simulation setting. Turning to the results, the mean biases in the third column of Table 1 are not close to zero reflecting that $\widehat{\mathrm{DID}}$ cannot match the ATE in the presence of misclassification. This is due to some actually treated groups (individuals) entering the comparison group that is supposed to be non treated. If treatment effects change over time, the inclusion of already treated groups in the reported pre-treatment periods is also an issue. As the fourth column shows, the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$ captures the ATE under constant treatment effects, but not under time-varying effects. Under constant treatment effects there are no wrong comparisons and the exact timing of the treatment does not matter since the effect size is the same in all periods. The mean bias of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is close to zero in all cases and decreases as the total number of groups increases.
Table 2 depicts the Root Mean Squared Error (RMSE). For the $\widehat{\mathrm{DID}}$ estimator, the RMSE is larger under constant treatment effects than under time varying treatment effects (see column 3), which reflects the same pattern seen for the mean bias. The RMSE of the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ is of similar magnitude or larger than those of the estimator for observed switchers $\widehat{\mathrm{DID}}^{S}$, even though the former do not exhibit a bias. This is to be expected since $\widehat{\mathrm{DID}}^{S^{*}}$ requires the estimation of unobserved switching probabilities. Not surprisingly, the estimator for true switchers $\widehat{\mathrm{DID}}^{S^{*}}$ produces a larger RMSE under time-varying treatment effects than under constant treatment effects given that the mean bias is close to zero in both cases.
# Groups DGP TE Mean Bias
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE -2.3332 0.0317 -0.0298
Time-vary. TE -2.0845 0.2536 -0.0432
G = 300 Constant TE -2.3287 -0.0029 -0.0020
Time-vary. TE -2.0786 0.2139 -0.0190
G = 600 Constant TE -2.3307 -0.0136 -0.0010
Time-vary. TE -2.0818 0.1981 -0.0098
Table 1: Results of 2,000 Monte Carlo replications for 100, 300 and 600 groups are presented. Columns 3-5 contain the simulation means of the difference between the estimates and the mean treatment effects. Column 3 refers to the De Chaisemartin and D’Haultefoeuille (2020) estimator $\widehat{\mathrm{DID}}$, column 4 to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ and column 5 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$.
# Groups DGP TE RMSE
$\bf{\widehat{\mathrm{DID}}}$ $\bf{\widehat{\mathrm{DID}}^{S}}$ $\bf{\widehat{\mathrm{DID}}^{S^{*}}}$
G = 100 Constant TE 2.4241 0.2050 0.2312
Time-vary. TE 2.1731 0.3835 0.8699
G = 300 Constant TE 2.4080 0.1119 0.1102
Time-vary. TE 2.1513 0.2713 0.2842
G = 600 Constant TE 2.4072 0.0820 0.0783
Time-vary. TE 2.1514 0.2312 0.1197
Table 2: Results of 2,000 Monte Carlo replications for 100, 300 and 600 groups are presented. Columns 3-5 contain the root mean squared errors (RMSE) between the estimates and the mean treatment effects. Column 3 refers to the De Chaisemartin and D’Haultefoeuille (2020) estimator $\widehat{\mathrm{DID}}$, column 4 to the DID estimator of observed switchers $\widehat{\mathrm{DID}}^{S}$ and column 5 to the DID estimator of true switchers $\widehat{\mathrm{DID}}^{S^{*}}$.
Next, we turn to the specification tests from Section 5. We make a few minor adjustments to the simulation framework presented above. The number of time periods is decreased to $T=10$ since the asymptotics for $T^{m}_{1,0}$ and $T^{s}_{1,0}$ rely on a small to moderate number of elements $\hat{\tau}_{1,0}(t,l)$ relative to the sample size $G$. Specifically, $\hat{\tau}_{0,0}(t,l)$ and $\hat{\tau}_{1,0}(t,l)$ are defined for $t\in\{(l+1),\ldots,T\}$ and $l\in\{2,\ldots,T-1\}$. This results in a cardinality of $|\mathcal{C}_{S}|=28$, i.e. 28 elements enter the statistics $T^{m}_{1,0}$ and $T^{s}_{1,0}$. The specification tests are designed to detect a violation of parallel trends in Assumption 5 and misclassification. Therefore, the data generating process entails 4 scenarios now:
1. Parallel Trends in pre-treatment periods and misclassification (PTM)
2. Trend violation in pre-treatment periods and misclassification (TVM)
3. Parallel Trends in pre-treatment periods and no misclassification (PTN)
4. Trend violation in pre-treatment periods and no misclassification (TVN)
While misclassification is simulated as before, a violation of the PT assumption in pre-treatment periods is simulated by adding a normalized time and group dependent trend: $Y_{g,t}^{TV}(D_{g,t}^{*})=Y_{g,t}(D_{g,t}^{*})+\frac{t\cdot g}{G}$. For simplicity, the treatment effect is kept constant over time now. The bootstrap is carried out as described in Section 5, and we use the 95-th percentile of the bootstrap distribution as critical value. To decrease the computing time we make use of the warp bootstrap method (Giacomini et al., 2013) and only draw a single bootstrap subsample within each Monte Carlo replication, resulting in 2,000 bootstrap samples.
Table 3 summarizes the rejection rates of the 4 tests for the different DGPs and group sizes. Column 3 is based on PT and misclassification (PTM), meaning that the rejection rates for the PT test in pre-treatment periods should converge to 0.05, while for the MC test we ought to observe a convergence towards one. Examining the results, for $G=100$, we see that the PT and MC tests underreject relative to the nominal level, even though the underrejections become smaller as $G$ grows. That said, even for $G=600$, the PT tests still slightly underreject in the presence of parallel trends (both with and without misclassification). Power results are generally favorable even for sample sizes as small as $G=100$. Note that the MC test rejects in the absence of misclassification and in the presence of parallel trends, since both patterns are not distinguishable when parallel trends are violated.
# Groups Tests $\textbf{H}_{\text{PTM}}$ $\textbf{H}_{\text{TVM}}$ $\textbf{H}_{\text{PTN}}$ $\textbf{H}_{\text{TVN}}$
G = 100 Sum PT 0.0300 0.9235 0.0295 0.8345
Max PT 0.0155 0.9215 0.0145 0.8015
Sum MC 0.9225 0.8945 0.0360 0.9145
Max MC 0.6655 0.6010 0.0210 0.8975
G = 300 Sum PT 0.0355 1 0.0420 1
Max PT 0.0300 1 0.0475 1
Sum MC 0.9960 1 0.0320 1
Max MC 0.9795 1 0.0250 1
G = 600 Sum PT 0.0410 1 0.0450 1
Max PT 0.0375 1 0.0300 1
Sum MC 0.9985 1 0.0445 1
Max MC 0.9935 1 0.0400 1
Table 3: Rejection rates of the four tests for different group numbers and for the four scenarios: PTM (parallel trends and misclassification), TVM (trend violation and misclassification), PTN (parallel trends and no misclassification) and TVN (trend violation and no misclassification).
7 Empirical Illustration
To illustrate the proposed estimators for true and observed switchers as well as the testing procedure, we revisit the application of Berkhout et al. (2024). The authors analyze how the introduction of a computer based testing system designed to curb cheating in high stakes national school examinations affected exam scores in Indonesia. As reported in their paper, cheating was a widespread problem in the nation-wide examinations at the end of junior and senior secondary school. Therefore, the Indonesian government gradually implemented computer based testing (CBT) instead of paper based exams, which made cheating highly unlikely and led to a drop in exam scores. Berkhout et al. (2024) point to anticipatory behavior by schools aware of their future treatment status, as treatment follows voluntary sign-up for CBT. That is, in order to soften the drop in exam scores, schools might try to counteract cheating practices prior to their actual treatment. The empirical findings of Berkhout et al. (2024) support the presence of such anticipatory behavior one year before the CBT implementation.
For the empirical analysis, we follow Berkhout et al. (2024) and use a balanced panel data set on exam results between 2010 and 2019. CBT was introduced between 2015-2019, but the 2015 and 2016 cohorts have been excluded, since their characteristics are different from the comparison groups. The number of observations is thus 38,621 schools for 10 years, see the original paper for further details on the data.
Cohort Estimators
Original $\widehat{\mathrm{DID}}^{S}_{t}$ $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ $\widehat{\lambda}_{t-1}$
Cohort 2017 $-5.21^{\ast}$ $-4.63^{\ast}$ / /
Cohort 2018 $-7.45^{\ast}$ $-7.50^{\ast}$ $-7.83^{\ast}$ 0.1580.1580.158
Cohort 2019 $-5.69^{\ast}$ $-5.69^{\ast}$ $-5.91^{\ast}$ 0.2970.2970.297
Combined $-6.30^{\ast}$ $-6.16^{\ast}$ $-7.00^{\ast}$ /
Table 4: Comparison of estimators for ATEs in the cohorts 2017, 2018, 2019 and a weighted combination of the cohorts. The star reflects significance at the 5% level.
Table 4 summarizes the treatment effect estimates from the original paper of Berkhout et al. (2024, Figure 5 and Table A3) and from our two proposed estimators as well as the estimated anticipation probabilities. Berkhout et al. (2024) use the staggered adoption estimator of Callaway and Sant’Anna (2021) where never treated schools serve as the control group and one year of anticipation is allowed for. We report both the estimates of the cohort specific effects and of the average effects weighted by the number of groups in each cohort. All estimated ATEs are statistically significant at the 5% level. [^13] We construct critical values for $H_{0}:\;ATE=0$ against its negation using the bootstrap as described in Section 5 with $B=999$. The estimates from the ATE estimator for observed switchers are very similar to the ATE estimates in the original study. Conceptually this is to be expected since the estimator of Callaway and Sant’Anna (2021) should also recover the ATE of observed switchers as the bias from anticipation in the treatment group is offset by dropping the first pre-treatment time period, and the bias from anticipation in the comparison group is avoided by using never treated groups as a comparison. In fact, the small numerical differences between our estimator and the estimator of Callaway and Sant’Anna (2021) are likely to stem from the larger comparison group used in our estimator $\widehat{\mathrm{DID}}^{S}_{t}$.
The cohort specific estimates of the ATE of True Switchers from $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ on the other hand are somewhat larger in absolute terms than in the original study. [^14] Note that the estimate for $\widehat{\mathrm{DID}}^{S^{\ast}}_{2017}$ is missing, since an already treated comparison group is lacking. Specifically, we see that for the 2018 cohort there is a drop in mean exam scores by 7.83 points, while for the 2019 cohort mean scores are estimated to drop by 5.91 points (scores range from 0 to 100). These somewhat larger decreases relative to the first and second column of Table 4 could capture not only the effect of CBT, but also the effect of anti-cheating measures implemented by schools in anticipation of the start of CBT. Moreover, since the size of the estimated treatment effects decline in magnitude with the post-treatment time periods as reported in Berkhout et al. (2024), it is reasonable that $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ are larger in magnitude as they reflect the (average) treatment effect at the time when schools actually switched behavior. Note that the comparison between the combined estimates of the three estimators should be interpreted with caution, since the first two are weighted averages of three cohorts, while the third (estimator for ATE of observed switchers) is a weighted average of two cohorts which display larger treatment effects than the first cohort. Finally, we estimate the anticipation probability to be 15.8% for the 2018 cohort, and 29.7% for the 2019 cohort.
Next and for illustration purposes, the testing framework from Section 5 is applied to test for the presence of parallel pre-trends and anticipatory behavior. We start by carrying out the the tests using never treated schools are used as control groups similar to Berkhout et al. (2024). Moreover, tests are conducted separately for the three treatment cohorts 2017, 2018 and 2019, where we use 499 bootstrap draws to obtain (5%) critical values. Table 5 displays the test results for the null hypotheses of parallel pre-trends, $H_{0}^{PT}$, and of no misclassification (no anticipation), $H_{0}^{MC}$, for each cohort using one comparison time period. [^15] For instance, the PT test for the 2017 cohort compares the outcomes of schools that are treated in 2017 to schools that are never treated by differencing between their 2015 and 2014 outcomes. The 2016 outcome is discarded since it could be affected by anticipatory behavior. The misclassification test then compares the same schools and their outcomes in 2015 and 2016.
In line with Berkhout et al. (2024), $H_{0}^{PT}$ is rejected at the 5% level for the 2017 cohort and not rejected for the 2018 cohort. However, while Berkhout et al. (2024) detect no violation of parallel pre-trends for the 2019 cohort when comparing 2016 and 2017 outcomes, our test rejects the null here $H_{0}^{PT}$ suggesting a possible violation of parallel pre-trends also in that cohort. Similarly, the misclassification tests based on both the sum and the max statistic reject $H_{0}^{MC}$ in the pre-treatment time period for all three cohorts, while the results in Berkhout et al. (2024) only indicated anticipatory effects for the 2018 and 2019 cohorts only. All our test results remain unchanged when using 1% or 10% significance levels instead. Finally, unreported results also show that constructing the test statistics using not yet treated (rather than just never treated) schools as comparison groups lead to the same rejection decisions with one exception: The PT test for the 2018 cohort rejects at the 5% level, but does not reject at the 1% level. This suggests that the PT assumption might be fragile, at least in periods just prior to the anticipation period.
Cohort $\mathbf{H_{0}^{PT}}$ $\mathbf{H_{0}^{MC}}$
Sum PT Max PT Sum MC Max MC
Cohort 2017 Statistic 389503.5 624.1 59139.8 243.2
Critical Val. 3966.5 63.0 4079.1 63.9
Cohort 2018 Statistic 1737.1 41.7 131859.8 363.1
Critical Val. 3553.6 59.6 2680.8 51.8
Cohort 2019 Statistic 9536.7 97.7 140575.9 374.9
Critical Val. 2712.9 52.1 3134.7 56.0
Table 5: Tests for parallel pre-trends ($H_{0}^{PT}$) and no misclassification ($H_{0}^{MC}$) for cohorts 2017, 2018, and 2019 using the sum and max statistics.
8 Conclusion
This paper investigates the consequences of treatment misclassification and anticipation in staggered adoption Difference-in-Differences designs, showing that these forms of misspecification can induce bias in commonly used estimators such as the TWFE estimator and other estimators suitable for treatment effect heterogeneity. We characterize the structure of these biases, and propose bias-corrected estimators that recover well-defined causal parameters under misclassification and anticipation, namely the ATE for observed and actual switching-to-treatment units.
Additionally, we introduce a two-step testing procedure based on moment equality tests to detect the timing and extent of misspecification, providing a diagnostic framework for empirical researchers to assess the validity of DiD designs under staggered adoption.
To assess the finite sample performance of our estimators and the moment equality tests, we conduct a small scale Monte Carlo study. Our simulation results illustrate the favorable finite-sample performance of both the proposed estimators and the tests, confirming their robustness under misspecification.
[…]
9 Appendix
9.1 Appendix: Additional Quantities
To study the asymptotic behavior of the estimator $\widehat{\mathrm{DID}}_{t}^{S^{\ast}}$ in (11), we require some additional notation. To this end, paralleling the notation in De Chaisemartin and D’Haultefoeuille (2020), for $(d,d^{\prime},t)\in\{0,1\}\times\{0,1\}\times\{2,\cdots,T\}$, let
$P_{d,d^{\prime},t}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=d,D_{g,t-1}=d^{\prime}\}N_{g,t},$
$Q_{d,d^{\prime},t}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=d,D_{g,t-1}=d^{\prime}\}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right).$
We also define:
$P_{d,d^{\prime},t-1}^{[1]}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=d,D_{g,t-1}=d^{\prime}\}N_{g,t-1},$
$Q_{d,d^{\prime},t}^{*}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=d,D_{g,t-1}=d^{\prime}\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-2}\right),$
$Q_{d,d^{\prime},t}^{**}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t+1}=d,D_{g,t}=d^{\prime}\}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right),$
$Q_{d,d^{\prime},t-1}^{***}$ $:=\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=d,D_{g,t-1}=d^{\prime}\}N_{g,t-1}\left(Y_{g,t}-Y_{g,t-1}\right),$
where we set $Q_{d,d^{\prime},2}^{*}=0$ and $Q_{d,d^{\prime},T}^{**}=0$. We collect the following quantities into a vector, which will be relevant for later use:
$\mathbf{V}:=\left(\mathbf{V}_{2}^{\prime},\cdots,\mathbf{V}_{T}^{\prime}\right)^{\prime},$ (20)
where
$\mathbf{V}_{t}$ $:=\left(P_{1,0,t},P_{0,0,t},P^{[1]}_{1,0,t-1},P^{[1]}_{0,0,t-1},P^{[1]}_{1,1,t-1},Q_{1,0,t},Q_{0,0,t},Q^{*}_{1,0,t},Q^{*}_{0,0,t},Q^{**}_{1,0,t},Q^{**}_{0,0,t},Q^{***}_{1,0,t-1},Q^{***}_{1,1,t-1},\right)^{\prime}.$
and denote its asymptotic variance by:
$\mathbf{\Sigma}_{\mathbf{V}}:=\lim_{G\rightarrow\infty}\text{Var}\left[\sqrt{G}\ \mathbf{V}\right],$ (21)
For the specification test, we define a similar set of quantities. Specifically, let
$Q_{0,0}(t,l)\equiv\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=0,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right),$
$P_{0,0}(t,l)\equiv\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=0,D_{g,t-1}=0\}N_{g,t}.$
Accordingly, define:
$Q_{1,0}(t,l)\equiv\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right),$
as well as
$P_{1,0}(t,l)\equiv\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}.$
Next, we let:
$\mathbf{U}(t,l)=\left(\begin{array}[]{c}P_{0,0}(t,l)\\ Q_{0,0}(t,l)\\ P_{1,0}(t,l)\\ Q_{1,0}(t,l)\end{array}\right),$
[…]
$\mathbf{\Sigma}_{\mathbf{U}}\equiv\lim_{G\rightarrow\infty}\text{Var}\left[\sqrt{G}\ \mathbf{U}\right],$ (22)
a $(4\cdot|\mathcal{C}_{S}|)\times(4\cdot|\mathcal{C}_{S}|)$ matrix given by:
$\boldsymbol{\Sigma}_{\mathbf{U}}\equiv\left(\begin{array}[]{ccc}\boldsymbol{\Sigma}_{(t,l)}&\ldots&\boldsymbol{\Sigma}_{(t,l),(j,k)}\\ \vdots&\ddots&\vdots\\ \boldsymbol{\Sigma}_{(j,k),(t,l)}&\ldots&\boldsymbol{\Sigma}_{(j,k)}\end{array}\right)$
with elements:
$\boldsymbol{\Sigma}_{(t,l)}$
$=$ $\left(\begin{array}[]{cccc}\text{Var}\left[P_{0,0}(t,l)\right]&\text{Cov}\left[P_{0,0}(t,l),Q_{0,0}(t,l)\right]&0&0\\ \text{Cov}\left[P_{0,0}(t,l),Q_{0,0}(t,l)\right]&\text{Var}\left[Q_{0,0}(t,l)\right]&0&0\\ 0&0&\text{Var}\left[P_{1,0}(t,l)\right]&\text{Cov}\left[P_{1,0}(t,l),Q_{1,0}(t,l)\right]\\ 0&0&\text{Cov}\left[P_{1,0}(t,l),Q_{1,0}(t,l)\right]&\text{Var}\left[Q_{1,0}(t,l)\right]\end{array}\right)$
9.2 Appendix: Proofs
Proof of Proposition 1.
First, for Assumption 3, we have:
$\quad\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}\right]-\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}\right]$
$=\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}^{*},\mathbf{D}\right]\mid\mathbf{D}\right]-\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*},\mathbf{D}_{g}\right]\mid\mathbf{D}_{g}\right]$
$=\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}^{*}\right]\mid\mathbf{D}\right]-\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}_{g}\right]$
$=\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}\right]-\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}_{g}\right]$
$=\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}_{g}\right]-\mathrm{E}\left[\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}_{g}\right]$
=0=0,= 0 ,
where the second equality follows from condition (4), the third from (1), and the fourth from (5).
Next, for Assumption 4, we similarly get:
$\quad\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}\right]-\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]$
$=\mathrm{E}\left[\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*},\mathbf{D}_{g}\right]\mid\mathbf{D}_{g}\right]-\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]$
$=\mathrm{E}\left[\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\mid\mathbf{D}_{g}^{*}\right]\mid\mathbf{D}_{g}\right]-\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]$
$=\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]-\mathrm{E}\left[{Y}_{g,t}(0)-{Y}_{g,t-1}(0)\right]$
=0=0,= 0 ,
where the second equality follows from (3) and the third from (2). ∎
Proof of Proposition 2.
By Assumption 2 and the Frisch–Waugh–Lovell theorem, we have:
$\mathrm{E}\!\left[\widehat{\beta}_{fe}\mid\mathbf{D}\right]=\frac{\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\mathrm{E}\!\left[Y_{g,t}\mid\mathbf{D}\right]}{\sum_{(g,t)}N_{g,t}D_{g,t}\varepsilon_{g,t}},$
where $\varepsilon_{g,t}$ is the estimated residual from the unit-level regression of $D_{g,t}$ on group and time fixed effects:
$D_{g,t}=\alpha+\gamma_{g}+\lambda_{t}+\varepsilon_{g,t}.$
As remarked in the proof of Theorem 1 in De Chaisemartin and D’Haultefoeuille (2020), it holds by construction that:
$\sum_{t=1}^{T}N_{g,t}\varepsilon_{g,t}=0$ (23)
for all $g\in\{1,\ldots,G\}$ and
$\sum_{g=1}^{G}N_{g,t}\varepsilon_{g,t}=0$ (24)
for all $t\in\{1,\ldots,T\}$. Focusing on the numerator, we apply (23) and (24) to obatin:
$\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\mathrm{E}\!\left[Y_{g,t}\mid\mathbf{D}\right]$ $=\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\Bigl{(}\mathrm{E}[Y_{g,t}\mid\mathbf{D}]-\mathrm{E}[Y_{g,1}\mid\mathbf{D}]-\mathrm{E}[Y_{1,t}\mid\mathbf{D}]+\mathrm{E}[Y_{1,1}\mid\mathbf{D}]\Bigr{)}.$
For each individual term, by $Y_{i,g,t}=Y_{i,g,t}(0)+D_{i,g,t}^{*}\Delta$ and Assumption 3, we have:
$\mathrm{E}\left[Y_{g,t}\mid\mathbf{D}\right]$ $=\mathrm{E}\left[\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}Y_{i,g,t}(0)+D_{i,g,t}^{\ast}\Delta\mid\mathbf{D}\right]=\mathrm{E}\left[Y_{g,t}(0)\mid\mathbf{D}_{g}\right]+\Delta\left(\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{Pr}\left(D_{i,g,t}^{\ast}=1\mid\mathbf{D}\right)\right).$
for all $(g,t)$. Thus, by Assumptions 4 and 5, we have:
$\quad\ \sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\left(\mathrm{E}\left[Y_{g,t}\mid\mathbf{D}\right]-\mathrm{E}\left[Y_{g,1}\mid\mathbf{D}\right]-\mathrm{E}\left[Y_{1,t}\mid\mathbf{D}\right]+\mathrm{E}\left[Y_{1,1}\mid\mathbf{D}\right]\right)$
$=\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\left(\mathrm{E}\left[Y_{g,t}(0)\right]-\mathrm{E}\left[Y_{g,1}(0)\right]-\mathrm{E}\left[Y_{1,t}(0)\right]+\mathrm{E}\left[Y_{1,1}(0)\right]\right)$
$\quad+\Delta\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\left(\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{Pr}\left(D_{i,g,t}^{\ast}=1\mid\mathbf{D}\right)-\frac{1}{N_{g,1}}\sum_{i=1}^{N_{g,1}}\mathrm{Pr}\left(D_{i,g,1}^{\ast}=1\mid\mathbf{D}\right)\right.$
$\left.\qquad\qquad\qquad\qquad\quad-\frac{1}{N_{1,t}}\sum_{i=1}^{N_{1,t}}\mathrm{Pr}\left(D_{i,1,t}^{\ast}=1\mid\mathbf{D}\right)+\frac{1}{N_{1,1}}\sum_{i=1}^{N_{1,1}}\mathrm{Pr}\left(D_{i,1,1}^{\ast}=1\mid\mathbf{D}\right)\right)$
$=$ $\Delta\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\left(\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{Pr}\left(D_{i,g,t}^{\ast}=1\mid\mathbf{D}\right)\right),$
where the last line follows again from (23) and (24). Thus, we obtain:
$\mathrm{E}\left[\widehat{\beta}_{fe}\mid\mathbf{D}\right]=\Delta\frac{\sum_{(g,t)}N_{g,t}\varepsilon_{g,t}\left(\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{Pr}\left(D_{i,g,t}^{\ast}=1\mid\mathbf{D}\right)\right)}{\sum_{(g,t)}N_{g,t}D_{g,t}\varepsilon_{g,t}},$
and the result of the proposition immediately follows.
Proof of Proposition 3.
First, note that
$\mathrm{E}\left[\widehat{\operatorname{DID}}\right]$ $=\mathrm{E}\left[\sum_{t=2}^{T}\frac{N_{1,0,t}}{N_{S}}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}\mid\mathbf{D}\right]\right].$
Besides, by Assumptions 1-5:
[…]
Let $t\geq 2$. Suppose that there exists some switching-to-treatment group $g\in\{1,\cdots,G\}$ such that $D_{g,t-1}=0$, $D_{g,t}=1$. By Assumption 8, there exists also some comparison group $g^{\prime}\in\{1,\ldots,G\}$ with $g^{\prime}\neq g$ such that $D_{g^{\prime},t-1}=D_{g^{\prime},t}=0.$ In this case,
$\quad\ N_{1,0,t}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}\mid\mathbf{D}\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{D}\right]-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{D}\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad-\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad+\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$\quad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad+\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\mid\mathbf{D}\right]$
$\quad-\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right],$
where the second equality follows from the formula for $\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{{D}}\right]$, and the last from Assumption 7. If there does not exist any switching-to-treatment group $g\in\{1,\cdots,G\}$ such that $D_{g,t-1}=0$, $D_{g,t}=1$, the above formula still holds, as $\widehat{\operatorname{DID}}_{t}=0$, $\mathbb{I}\left\{D_{g,t}=1,D_{g,t-1}=0\right\}=0$ and $N_{1,0,t}=0$ in this case. Thus,
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ $=\mathrm{E}\left[\mathrm{E}\left[\widehat{\mathrm{DID}}\mid\mathbf{D}\right]\right]$
$=\mathrm{E}\left[\sum_{t\geq 2}\frac{N_{1,0,t}}{N_{S}}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}\mid\mathbf{D}\right]\right]$
$=\mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}(Y_{i,g,t}(1)-Y_{i,g,t}(0))\right]$
$\quad-\mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad-\mathrm{E}\left[\sum_{t\geq 2}\frac{N_{1,0,t}}{N_{S}N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$=\underbrace{\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}(Y_{i,g,t}(1)-Y_{i,g,t}(0))\right]}_{\delta^{S}}$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{1,0,t}}{N_{0,0,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
Proof of Proposition 4.
Recall from the previous proof that under Assumption 1-8:
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]=$ $\mathrm{E}\left[\mathrm{E}\left[\widehat{\mathrm{DID}}\mid\mathbf{D}\right]\right]$
$=$ $\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}(Y_{i,g,t}(1)-Y_{i,g,t}(0))\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{1,0,t}}{N_{0,0,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\equiv$ $A_{N_{S}}+B_{N_{S}}+C_{N_{S}}$
with the obvious notation. We only decompose $A_{N_{S}}$ and $C_{N_{S}}$, while $B_{N_{S}}$ will contribute directly to the bias term in the proposition. Now, noting that $\mathbb{I}\left\{D_{g,t}=1,D_{g,t-1}=0\right\}=\mathbb{I}\left\{D_{g,t}=1,D_{g,t-1}=0\right\}D_{g,t}^{*}$ by Assumption 7, and $D_{i,g,t}^{\ast}=S_{i,g,t}^{\ast}D_{i,g,t}^{\ast}+(1-S_{i,g,t}^{\ast})D_{i,g,t}^{\ast}$, we have for $A_{N_{S}}$ that:
$A_{N_{S}}$ $=\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D_{i,g,t}^{\ast}(Y_{i,g,t}(1)-Y_{i,g,t}(0))\right]$
$=\mathrm{E}\left[\frac{1}{N_{S^{\ast}}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$
On the other hand, the term $C_{N_{S}}$ can be decomposed as follows:
$C_{N_{S}}$ $=\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\left(\frac{N_{1,0,t}}{N_{0,0,t}}-1\right){D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\equiv C_{N_{S}}(1)+C_{N_{S}}(2),$
where $C_{N_{S}}(1)$ can be decomposed further as:
$C_{N_{S}}(1)$ $=\mathrm{E}\left[\frac{1}{N_{S^{\ast}}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$
Note in fact that the first term on the RHS of the last equation is equal to:
$\mathrm{E}\left[\frac{1}{N_{S^{\ast}}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-1}}{D}_{i,g,t-1}^{*}S_{i,g,t-1}^{\ast}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
by the assumptions on the switching probabilities in Proposition 4. Finally, we have that:
$\delta^{S^{*}}$ $=\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(i,g,t):t\geq 2,D_{i,g,t}^{\ast}=1,S_{i,g,t}^{*}=1}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$=\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(i,g,t):t\geq 2,D_{i,g,t}^{\ast}=1,S_{i,g,t}^{*}=1,D_{g,t}=1,D_{g,t-1}=0}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(i,g,t):t\geq 2,D_{i,g,t-1}^{\ast}=1,S_{i,g,t-1}^{*}=1,D_{g,t}=1,D_{g,t-1}=0}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right]$
$=\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D_{i,g,t}^{\ast}S_{i,g,t}^{*}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S^{*}}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-1}}D_{i,g,t-1}^{\ast}S_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right],$
where the second equality follows from the assumptions on the switching probabilities. Thus, collecting all terms, we obtain:
$\mathrm{E}\left[\widehat{\mathrm{DID}}\right]$ $=\mathrm{E}\left[\mathrm{E}\left[\widehat{\mathrm{DID}}\mid\mathbf{D}\right]\right]$
$=\delta^{S^{\ast}}$
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}D^{\ast}_{i,g,t}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad+\mathrm{E}\left[\left(\frac{1}{N_{S}}-\frac{1}{N_{S^{\ast}}}\right)\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}S_{i,g,t}^{\ast}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(1-S_{i,g,t}^{\ast}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad-\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(g,t):t\geq 2,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\left(\frac{N_{1,0,t}}{N_{0,0,t}}-1\right){D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right].$
Proof of Proposition 5.
First, we consider $\widehat{\operatorname{DID}}_{t}^{**}$. For $2\leq t\leq T-1$, suppose that there exists a switching-to-treatment group $g\in\{1,\cdots,G\}$ such that $D_{g,t}=0$, $D_{g,t+1}=1$, then, by Assumption 8, there exists also some comparison group $g^{\prime}\in\{1,\ldots,G\}$ with $g^{\prime}\neq g$ such that $D_{g^{\prime},t}=D_{g^{\prime},t+1}=0$. As a consequence of Assumption 6, $D_{g^{\prime},t-1}=0$, and Assumption 1 implies that all denominators are nonzero. Recall that $N_{1,0,t}^{[1]}:=\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}$ and $N_{0,0,t}^{[1]}:=\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}$. By (25) in the proof of Proposition 3:
$\quad\ \frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{**}\mid\mathbf{D}\right]$
$=\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{D}\right]-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{D}\right]$
$=\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad+\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$\quad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$=\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
where the second equality follows from the formula for $\mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{{D}}\right]$, and the last from Assumptions 6 and 7. If there exists no switching-to-treatment group $g\in\{1,\cdots,G\}$ such that $D_{g,t}=0$, $D_{g,t+1}=1$, the above formula still holds, as $\widehat{\operatorname{DID}}_{t}^{**}=0$ and $\mathbb{I}\left\{D_{g,t+1}=1,D_{g,t}=0\right\}=0$. Note that in this case, $N_{0,0,t}$ must be nonzero, unless there is no group observed to switch at any time at all (in which case we can set all $\frac{1}{N_{0,0,t}}=0$). This holds true because otherwise, by Assumption 6, it must be $D_{g,t}=1,D_{g,t-1}=0$ or $D_{g,t}=1,D_{g,t-1}=1$ for all $g\in\{1,\cdots,G\}$, and one can show either case violates Assumption 8. The last term can be further written as:
$\quad\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad+\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$=\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
When $t=T$, the formula still continues to hold as $\widehat{\operatorname{DID}}_{T}^{**}=0$ and the last term becomes zero under Assumption 9.
For $3\leq t\leq T$, similar steps yield:
$\quad\ N_{1,0,t}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{*}\mid\mathbf{D}\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[Y_{g,t-1}-Y_{g,t-2}\mid\mathbf{D}\right]$
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t+1}\mathrm{E}\left[Y_{g,t-1}-Y_{g,t-2}\mid\mathbf{D}\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]$
$\quad-\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\quad+\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]$
$\quad+\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
$=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]$
When $t=2$, the formula still continues to hold as $\widehat{\operatorname{DID}}_{2}^{*}=0$ and the last term becomes zero under Assumption 10.
Finally, combining the above, we obtain:
$\quad\ \mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\left(N_{1,0,t}\widehat{\operatorname{DID}}_{t}+N_{1,0,t}\widehat{\operatorname{DID}}_{t}^{*}+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\widehat{\operatorname{DID}}_{t}^{**}\right)\right]$
$=\mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}Y_{i,g,t}(1)-Y_{i,g,t}(0)\right]$
$\quad-\mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-1}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad-\mathrm{E}\left[\sum_{t\geq 2}\frac{N_{1,0,t}}{N_{S}N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$\quad+\mathrm{E}\left[\sum_{t\geq 2}\frac{1}{N_{S}}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-1}}\frac{N_{g,t}}{N_{g,t-1}}{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\right]$
$\quad+\mathrm{E}\left[\sum_{t\geq 2}\frac{N_{1,0,t}}{N_{S}N_{0,0,t}}\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t}}{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\right]$
$=\mathrm{E}\left[\frac{1}{N_{S}}\sum_{(i,g,t):t\geq 2,D_{g,t}=1,D_{g,t-1}=0}Y_{i,g,t}(1)-Y_{i,g,t}(0)\right].$
Proof of Proposition 6.
Recall the definition of $\mathbf{V}=\left(\mathbf{V}_{2}^{\prime},\cdots,\mathbf{V}_{T}^{\prime}\right)^{\prime}$ in (20) , where
$\mathbf{V}_{t}$ $=\left(P_{1,0,t},P_{0,0,t},P^{[1]}_{1,0,t-1},P^{[1]}_{0,0,t-1},P^{[1]}_{1,1,t-1},Q_{1,0,t},Q_{0,0,t},Q^{*}_{1,0,t},Q^{*}_{0,0,t},Q^{**}_{1,0,t},Q^{**}_{0,0,t},Q^{***}_{1,0,t-1},Q^{***}_{1,1,t-1},\right)^{\prime}.$
Thus, by some algebra, for all combinations of $P_{d,d^{\prime},t}$, $P_{d,d^{\prime},t-1}^{[1]}$ and $\widehat{\lambda}_{t-1}$, with $(d,d^{\prime},t)\in\{0,1\}\times\{0,1\}\times\{2,\cdots,T\}$ such that the denominator is strictly positive, we can write
$\quad\ \widehat{\mathrm{DID}}^{S^{\ast}}$
$=\frac{\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}\left(N_{1,0,t-1}^{[1]}\widehat{\operatorname{DID}}_{t-1}^{**}+N_{1,0,t-1}^{[1]}\widehat{\operatorname{DID}}_{t-1}^{***}\right)+\left(1-\widehat{\lambda}_{t-1}\right)\left(N_{1,0,t}\widehat{\operatorname{DID}}_{t}+N_{1,0,t}\widehat{\operatorname{DID}}_{t}^{*}+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\widehat{\operatorname{DID}}_{t}^{**}\right)\right)}{\sum_{t\geq 2}\sum_{D_{g,t}=1,D_{g,t-1}=0}\left(\widehat{\lambda}_{t-1}N_{g,t-1}+(1-\widehat{\lambda}_{t-1})N_{g,t}\right)}$
$=\frac{\sum_{t\geq 2}\widehat{\lambda}_{t-1}\left(P_{1,0,t-1}^{[1]}\left(\frac{Q_{1,0,t-1}^{**}}{P_{1,0,t-1}^{[1]}}-\frac{Q_{0,0,t-1}^{**}}{P_{0,0,t-1}^{[1]}}\right)+P_{1,0,t-1}^{[1]}\left(\frac{Q_{1,0,t-1}^{***}}{P_{1,0,t-1}^{[1]}}-\frac{Q_{1,1,t-1}^{***}}{P_{1,1,t-1}^{[1]}}\right)\right)}{\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}P^{[1]}_{1,0,t-1}+(1-\widehat{\lambda}_{t-1})P_{1,0,t}\right)}$
$\quad+\frac{\sum_{t\geq 2}\left(1-\widehat{\lambda}_{t-1}\right)\left(P_{1,0,t}\left(\frac{Q_{1,0,t}}{P_{1,0,t}}-\frac{Q_{0,0,t}}{P_{0,0,t}}\right)+P_{1,0,t}\left(\frac{Q_{1,0,t}^{*}}{P_{1,0,t}}-\frac{Q_{0,0,t}^{*}}{P_{0,0,t}}\right)+\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}}\left(\frac{Q_{1,0,t}^{**}}{P_{1,0,t}^{[1]}}-\frac{Q_{0,0,t}^{**}}{P_{0,0,t}^{[1]}}\right)\right)}{\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}P^{[1]}_{1,0,t-1}+(1-\widehat{\lambda}_{t-1})P_{1,0,t}\right)}$
$=f_{N}\left(\mathbf{V}\right)\times f_{D}\left(\mathbf{V}\right)=:f\left(\mathbf{V}\right),$
where
$f_{N}\left(\mathbf{V}\right)$ $:=\sum_{t\geq 2}\left(g_{t}(\boldsymbol{V})\cdot\left(\left(Q_{1,0,t-1}^{**}-\frac{P_{1,0,t-1}^{[1]}}{P_{0,0,t-1}^{[1]}}Q_{0,0,t-1}^{**}\right)+\left(Q_{1,0,t-1}^{***}-\frac{P_{1,0,t-1}^{[1]}}{P_{1,1,t-1}^{[1]}}Q_{1,1,t-1}^{***}\right)\right)\right.$
$\qquad\qquad\left.+\left(1-g_{t}(\boldsymbol{V})\right)\cdot\left(\left(Q_{1,0,t}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}\right)+\left(Q_{1,0,t}^{*}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}^{*}\right)+\left(\frac{P_{1,0,t}}{P_{0,0,t}}Q_{1,0,t}^{**}-\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}^{[1]}P_{0,0,t}}Q_{0,0,t}^{**}\right)\right)\right)$
and
$f_{D}\left(\mathbf{V}\right):=\left(\sum_{t\geq 2}\left(g_{t}(\boldsymbol{V})P^{[1]}_{1,0,t-1}+(1-g_{t}(\boldsymbol{V}))P_{1,0,t}\right)\right)^{-1}$
with
$g_{t}(\boldsymbol{V}):=\frac{\displaystyle\frac{Q_{1,0,t-1}^{**}}{P_{1,0,t-1}^{[1]}}-\frac{Q_{0,0,t-1}^{**}}{P_{0,0,t-1}^{[1]}}}{\displaystyle\left(\frac{Q_{1,0,t-1}^{**}}{P_{1,0,t-1}^{[1]}}-\frac{Q_{0,0,t-1}^{**}}{P_{0,0,t-1}^{[1]}}\right)+\left(\frac{Q_{1,0,t-1}^{***}}{P_{1,0,t-1}^{[1]}}-\frac{Q_{1,1,t-1}^{***}}{P_{1,1,t-1}^{[1]}}\right)}=\widehat{\lambda}_{t-1}.$
Note that with the above definitions, we can write:
$f\left(\mathrm{E}\left[\mathbf{V}\right]\right)=f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\times f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right),$
where
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
$=\sum_{t\geq 2}\left(g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]\right)\right)\right.$
$\qquad\qquad+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$
and
$f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\left(\sum_{t\geq 2}\left(g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\mathrm{E}\left[P^{[1]}_{1,0,t-1}\right]+(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right))\mathrm{E}\left[P_{1,0,t}\right]\right)\right)^{-1}$
with
$g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\frac{\displaystyle\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}}{\displaystyle\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\right)+\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\right)}$
In the following, we will start by arguing that
$\sqrt{G}\left(f\left(\mathbf{V}\right)-f\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}\left(0,\sigma^{2}_{\delta^{S^{\ast}}}\right).$
In a second step, we will show that
$\sqrt{G}\left(f\left(\mathrm{E}\left[\mathbf{V}\right]\right)-\delta^{S^{\ast}}\right)=o(1),$
which together imply the desired result by Slutsky’s theorem.
Step 1: We begin by analyzing the components of $\mathbf{V}$ in (20), focusing on $Q_{1,0,t}=\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}(Y_{g,t}-Y_{g,t-1})$. The arguments for the remaining terms follow analogously. For a given period $t\geq 2$, we denote $Z_{g}:=\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t}-Y_{g,t-1}\right)$, $\mu_{g}:=\mathrm{E}\left[Z_{g}\right]$ and $\sigma_{g}^{2}:=\mathrm{Var}\left[Z_{g}\right]$. We observe:
$\sup_{g}\mathrm{E}\left[Z_{g}^{4}\right]$ $\leq\sup_{g}N_{g,t}^{4}\times\sup_{g}\mathrm{E}\left[(Y_{g,t}-Y_{g,t-1})^{4}\right]$
$\leq\sup_{g}N_{g,t}^{4}\times\left(8\sup_{g}\mathrm{E}\left[Y_{g,t}^{4}\right]+8\sup_{g}\mathrm{E}\left[Y_{g,t-1}^{4}\right]\right)$
$\leq\sup_{g}N_{g,t}^{4}\times 16\sup_{i,g}\mathrm{E}\left[Y_{i,g,t}^{4}\right]$
$\leq C_{1}$
for some constant $C_{1}>0$, where the second line follows from the elementary inequality $|a+b|^{p}\leq 2^{p-1}\left(|a|^{p}+|b|^{p}\right)$ for all $a,b\in\mathbb{R}$ and $p>1$, the third from the Jensen’s inequality to the convex function $x\mapsto x^{4}$, and the last from Assumption 18. Applying again the above elementary inequality and Jensen’s inequality to $x\mapsto x^{4}$, this implies a uniform bound over $g$:
$\mathrm{E}\left[\left|Z_{g}-\mu_{g}\right|^{4}\right]\leq 8\mathrm{E}\left[Z_{g}^{4}\right]+8\mathrm{E}\left[Z_{g}\right]^{4}\leq 16\mathrm{E}\left[Z_{g}^{4}\right]\leq 16C_{1}.$
Finally, since $\operatorname{Var}\left(\sqrt{G}Q_{1,0,t}\right)=\frac{1}{G}\sum_{g=1}^{G}\operatorname{Var}\left(Z_{g}\right)$ under Assumptions 14, Assumption 18 implies that
$\frac{S_{G}^{2}}{G}:=\frac{1}{G}\sum_{g=1}^{G}\sigma_{g}^{2}=\operatorname{Var}\left(\sqrt{G}Q_{1,0,t}\right)\xrightarrow{G\rightarrow\infty}C_{2}$
for some constant $C_{2}>0$, so that
$\frac{C_{2}}{2}\leq\frac{S_{G}^{2}}{G}$
for sufficiently large $G$ by the definition of sequence limit. Hence, we verify Lyapunov’s condition with $\delta=2$:
$\frac{1}{S_{G}^{2+\delta}}\sum_{g=1}^{G}\mathrm{E}\!\left[\lvert Z_{g}-\mu_{g}\rvert^{2+\delta}\right]\leq\frac{16C_{1}G}{\left(\frac{C_{2}}{2}G\right)^{2}}=\frac{64C_{1}}{C_{2}^{2}}\frac{1}{G}\xrightarrow{G\rightarrow\infty}0.$
Lyapunov’s CLT thus yields the scalar convergence
$\sqrt{G}\left(Q_{1,0,t}-\mathrm{E}\left[Q_{1,0,t}\right]\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,C_{2}\right).$
Applying the same argument to every coordinate of $\mathbf{V}$ and to any linear combination $\mathbf{c}^{\prime}\mathbf{V}$ (with $\mathbf{c}^{\prime}$ fixed and of suitable dimension), and invoking the Cramér–Wold device, we obtain
$\sqrt{G}\left(\mathbf{V}-\mathrm{E}\left[\mathbf{V}\right]\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{V}}),$
where $\boldsymbol{\Sigma}_{\mathbf{V}}=\lim_{G\rightarrow\infty}\text{Var}\left[\sqrt{G}\ \mathbf{V}\right]$ as in (21).
Now, by Assumption 18, the limit $\mathbf{V}^{\infty}:=\lim_{G\to\infty}\mathrm{E}[\mathbf{V}]$ exists. Besides, paralleling the existence argument in De Chaisemartin and D’Haultefoeuille (2020), one can show that $f$ is differentiable in a neighborhood of $\mathbf{V}^{\infty}$. Thus, the uniform Delta method (see, e.g., Van der Vaart, 1998, Theorem 3.8) yields:
$\sqrt{G}\left(f\left(\mathbf{V}\right)-f\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)$ $=\mathbf{J}_{f}\left(\mathbf{V}^{\infty}\right)\times\sqrt{G}\left(\mathbf{V}-\mathrm{E}\left[\mathbf{V}\right]\right)+o_{\Pr}(1)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}\left(0,\sigma^{2}_{\delta^{S^{\ast}}}\right),$
where $\sigma^{2}_{\delta^{S^{\ast}}}=\mathbf{J}_{f}\bigl{(}\mathbf{V}^{\infty}\bigr{)}\boldsymbol{\Sigma}_{\mathbf{V}}\mathbf{J}_{f}\bigl{(}\mathbf{V}^{\infty}\bigr{)}^{\prime}$.
Step 2.1: Next, we continue to show that
$\sqrt{G}\left(f\left(\mathrm{E}\left[\mathbf{V}\right]\right)-\delta^{S^{\ast}}\right)=o(1).$
To this end, we define
$T^{S}:=\frac{1}{G}\sum_{(i,g,t):t\geq 2,S_{i,g,t}^{*}=1}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\quad\text{ and }\quad P^{S}:=\frac{N_{S^{*}}}{G},$
so that
$\delta^{S^{*}}=\mathrm{E}\left[\frac{T^{S}}{P^{S}}\right].$
In the following, we will show $f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\mathrm{E}\left[T^{S}\right]$ and $f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\left(\mathrm{E}\left[P^{S}\right]\right)^{-1}$, so that
$f\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\frac{\mathrm{E}\left[T^{S}\right]}{\mathrm{E}\left[P^{S}\right]},$
where
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
$=\sum_{t\geq 2}\left(g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]\right)\right)\right.$
$\qquad\qquad+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$
and
$f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\left(\sum_{t\geq 2}\left(g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\mathrm{E}\left[P^{[1]}_{1,0,t-1}\right]+(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right))\mathrm{E}\left[P_{1,0,t}\right]\right)\right)^{-1}$
with
$g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\frac{\displaystyle\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}}{\displaystyle\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\right)+\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\right)}$
Then, since $\lim_{G\rightarrow\infty}\mathrm{E}\left[P^{S}\right]>0$ by Assumption 17, the problem reduces to showing
$\sqrt{G}\left(\mathrm{E}\left[T^{S}\right]-\mathrm{E}\left[P^{S}\right]\mathrm{E}\left[\frac{T^{S}}{P^{S}}\right]\right)=o(1).$
Examining the components in $f\left(\mathrm{E}\left[\mathbf{V}\right]\right)$, we first observe, by (25) in the proof of Proposition 3 (which continue to hold under Assumptions 1-2 and 14-16), that
$\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]$ $=\mathrm{E}\left[\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}(Y_{g,t-1}-Y_{g,t-2})\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\frac{N_{g,t-1}}{N_{g,t-2}}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}\right]\times\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}\right]\times\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right],$
where the the third equality follows from Assumption 7. By similar arguments:
$\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]$ $=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=0,D_{g,t-1}=0\}N_{g,t-1}\right]\times\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right],$
$\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]$ $=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}\right],$
$\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]$ $=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=0,D_{g,t-1}=0\}N_{g,t-1}\right].$
Combining the above, we obtain
$\quad\ \mathrm{E}\left[Q_{1,0,t-1}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\mid{D}_{i,g,t-1}^{*}=1,\mathbf{D}_{g}\right]\mathrm{Pr}\left[{D}_{i,g,t-1}^{\ast}=1\mid\mathbf{{D}}_{g}\right]\right]$
$=\lambda_{t-1}\times\frac{1}{G}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right]$
where the last step follows from Assumptions 12 and 13, and the fact that under Assumptions 6:
$\quad\ \mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\times\mathrm{Pr}\left[{D}_{i,g,t-1}^{\ast}=1\mid\mathbf{{D}}_{g}\right]$
$=\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\times\mathrm{Pr}\left[{D}_{i,g,t-1}^{\ast}=1\mid{D}_{g,t}=1,D_{g,t-1}=0\right]$
almost surely. Under the same set of assumptions, one can also show, in parallel to (25), that:
[…]
Using this equation, we obtain:
$\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]$ $=\mathrm{E}\left[\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}(Y_{g,t}-Y_{g,t-1})\right]$
$=-\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\frac{N_{g,t-1}}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[\left(1-D_{i,g,t}^{*}\right)\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[\left(1-D_{i,g,t-1}^{*}\right)\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\cdot N_{g,t}\right]\cdot\mathrm{E}\left[Y_{1,t}(1)-Y_{1,t-1}(1)\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[\left(1-{D}_{i,g,t-1}^{*}\right)\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$\quad+\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t-1}\right]\times\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right],$
where the the third equality follows from Assumption 7. Repeating the above arguments yield:
$\quad\ \mathrm{E}\left[Q_{1,0,t-1}^{***}\right]-\frac{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[\left(1-{D}_{i,g,t-1}^{*}\right)\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}_{g}\right]\right]$
$=\frac{1}{G}\mathrm{E}\left[\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\mid{D}_{i,g,t-1}^{*}=0,\mathbf{D}_{g}\right]\mathrm{Pr}\left[{D}_{i,g,t-1}^{\ast}=0\mid\mathbf{{D}}_{g}\right]\right]$
$=\left(1-\lambda_{t-1}\right)\times\frac{1}{G}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right]$
where the last line follows from Assumption 12. Therefore, we obtain
$g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\frac{\displaystyle\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}}{\displaystyle\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{0,0,t-1}^{**}\right]}{\mathrm{E}\left[P_{0,0,t-1}^{[1]}\right]}\right)+\left(\frac{\mathrm{E}\left[Q_{1,0,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,0,t-1}^{[1]}\right]}-\frac{\mathrm{E}\left[Q_{1,1,t-1}^{***}\right]}{\mathrm{E}\left[P_{1,1,t-1}^{[1]}\right]}\right)}=\lambda_{t-1}.$
Finally, since:
$\widehat{\mathrm{DID}}_{t}^{S}$
$=$ $\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)$
$+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)$
$=$ $\frac{1}{G}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right].$
we also obtain:
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
$=\mathrm{E}\left[\frac{1}{G}\sum_{t\geq 2}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\left(\lambda_{t-1}\times\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right.\right.\left.+\left(1-\lambda_{t-1}\right)\times\left.\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right)\right],$
and
$f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$ $=\left(\sum_{t\geq 2}\left(\lambda_{t-1}\mathrm{E}\left[P^{[1]}_{1,0,t-1}\right]+\left(1-\lambda_{t-1}\right)\mathrm{E}\left[P_{1,0,t}\right]\right)\right)^{-1}.$
On the other hand, note that by Bayes’ rule, one can show that in staggered adoption designs, Assumptions 11 implies that
$\mathrm{P}\!\left(D_{g,t}=1,D_{g,t-1}=0\mid S^{\ast}_{i,g,t}=1\right)+\mathrm{P}\!\left(D_{g,t+1}=1,D_{g,t}=0\mid S^{\ast}_{i,g,t}=1\right)=1,\quad t\in\{2,\cdots,T-1\},$
and similarly:
$\mathrm{P}\!\left(D_{g,T}=1,D_{g,T-1}=0\mid S^{\ast}_{i,g,T}=1\right)=1.$
Thus,
$\quad\ \mathrm{E}\left[P^{S}\right]=\mathrm{E}\left[\frac{N_{S^{*}}}{G}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S^{*}_{i,g,t}=1\}\right]$
$=\mathrm{E}\left[\frac{1}{G}\sum_{t=2}^{T-1}\sum_{g=1}^{G}\left(\mathbb{I}\{D_{g,t+1}=1,D_{g,t}=0\}+\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\right)\sum_{i=1}^{N_{g,t}}\mathbb{I}\{S^{*}_{i,g,t}=1\}\right]$
$\quad+\mathrm{E}\left[\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{D_{g,T}=1,D_{g,T-1}=0\}\sum_{i=1}^{N_{g,T}}\mathbb{I}\{S^{*}_{i,g,T}=1\}\right]$
$=\mathrm{E}\left[\frac{1}{G}\sum_{t=3}^{T}\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t-1}}\mathbb{I}\{S_{i,g,t-1}^{*}=1\}+\sum_{t=2}^{T}\sum_{g=1}^{G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\sum_{i=1}^{N_{g,t}}\mathbb{I}\{S_{i,g,t}^{*}=1\}\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\mathrm{E}\left[\sum_{i=1}^{N_{g,t-1}}\mathbb{I}\left\{S_{i,g,t-1}^{\ast}=1\right\}+\sum_{i=1}^{N_{g,t}}\mathbb{I}\left\{S_{i,g,t}^{\ast}=1\right\}\mid\mathbf{D}_{g}\right]\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\left(\lambda_{t-1}N_{g,t-1}+\left(1-\lambda_{t-1}\right)N_{g,t}\right)\right]$
$=\sum_{t\geq 2}\left(\lambda_{t-1}\mathrm{E}\left[P^{[1]}_{1,0,t-1}\right]+\left(1-\lambda_{t-1}\right)\mathrm{E}\left[P_{1,0,t}\right]\right)=\left(f_{D}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)^{-1},$
where the second equality follows from reindexing, and the third from Assumptions 10 and 11, the fourth from $\mathrm{Pr}\left(S^{\ast}_{g,1}=0\mid D_{g,2}=0,\ D_{g,1}=0\right)=1$, and the sixth from Assumption 12. Similarly, we obtain:
$\quad\ \mathrm{E}\left[T^{S}\right]=\mathrm{E}\left[\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\left(\mathrm{E}\left[\sum_{i=1}^{N_{g,t-1}}\mathbb{I}\left\{S_{i,g,t-1}^{\ast}=1\right\}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{D}_{g}\right]\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\mathbb{I}\left\{S_{i,g,t}^{\ast}=1\right\}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid\mathbf{D}_{g}\right]\right)\right]$
$=\frac{1}{G}\sum_{t\geq 2}\sum_{g=1}^{G}\mathrm{E}\left[\mathbb{I}\left\{D_{g,t}=1,\ D_{g,t-1}=0\right\}\left(\mathrm{E}\left[\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid S_{i,g,t-1}^{\ast}=1,\mathbf{D}_{g}\right]\times\Pr\left(S_{i,g,t-1}^{\ast}=1\mid\mathbf{D}_{g}\right)\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \left.\left.+\mathrm{E}\left[\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid S_{i,g,t}^{\ast}=1,\mathbf{D}_{g}\right]\times\Pr\left(S_{i,g,t}^{\ast}=1\mid\mathbf{D}_{g}\right)\right)\right]$
$=\mathrm{E}\left[\frac{1}{G}\sum_{t\geq 2}\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\left(\lambda_{t-1}\times\sum_{i=1}^{N_{g,t-1}}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\right.\right.\left.+\left(1-\lambda_{t-1}\right)\times\left.\sum_{i=1}^{N_{g,t}}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\right)\right],$
where the last step follows from Assumptions 12 and 13.
Thus, we conclude that
$f\left(\mathrm{E}\left[\mathbf{V}\right]\right)=\frac{\mathrm{E}\left[T^{S}\right]}{\mathrm{E}\left[P^{S}\right]},$
and the problem reduces to showing $\sqrt{G}\left(\mathrm{E}\left[T^{S}\right]-\mathrm{E}\left[P^{S}\right]\mathrm{E}\left[\frac{T^{S}}{P^{S}}\right]\right)=o(1).$
Step 2.2: Next, we decompose:
$\mathrm{E}\left[T^{S}\right]-\mathrm{E}\left[P^{S}\right]\cdot\mathrm{E}\left[\frac{T^{S}}{P^{S}}\right]$ $=\underbrace{\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{\widetilde{T}^{S}}{P^{S}}\cdot I_{G}\right]}_{=:A}+\underbrace{\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{\widetilde{T}^{S}}{P^{S}}\cdot(1-I_{G})\right]}_{=:B},$ (26)
where
$\widetilde{T}^{S}=\mathrm{E}\left[T^{S}\mid\left(S_{i,g,t}^{\ast}\right)_{(i,g,t)\in\{1,\cdots,N_{g,t}\}\times\{1,\cdots,G\}\times\{1,\cdots,T\}}\right]\quad\text{and}\quad I_{G}:=\mathbb{I}\left\{\left|P^{S}-\mathrm{E}\left[P^{S}\right]\right|<\varepsilon_{G}\right\}$
for some $\varepsilon_{G}>0$ that will be specified below.
We begin with the second term $B$ on the right-hand side of (26). Applying the Cauchy–Schwarz inequality twice yields:
$\sqrt{G}\left|B\right|$ $=\sqrt{G}\left|\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{\widetilde{T}^{S}}{P^{S}}\cdot(1-I_{G})\right]\right|$
$\leq\sqrt{G}\cdot\left(\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)^{2}\right]\right)^{1/2}\cdot\left(\mathrm{E}\left[\left(\frac{\widetilde{T}^{S}}{P^{S}}\right)^{4}\right]^{1/2}\cdot\mathrm{E}\left[(1-I_{G})^{4}\right]^{1/2}\right)^{1/2}$
$=\underbrace{\sqrt{G}\cdot\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}_{=:B_{1}}\cdot\underbrace{\left(\mathrm{E}\left[\left(\frac{\widetilde{T}^{S}}{P^{S}}\right)^{4}\right]\cdot\mathrm{Pr}(I_{G}=0)\right)^{1/4}}_{=:B_{2}}.$
It is easy to check that the term $B_{1}$ converges to a finite limit by Assumption 18. It remains to show that the remaining term tends to zero. We observe that
$\widetilde{T}^{S}$ $=\mathrm{E}\left[T^{S}\mid\left(S_{i,g,t}^{\ast}\right)\right]=\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\cdot\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\mid\left(S_{i,g,t}^{\ast}\right)\right],$
$P^{S}$ $=\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}.$
Hence, for any fixed $G$, applying Jensen’s inequality to the convex function $x\mapsto x^{4}$ and the elementary inequality $|a+b|^{p}\leq 2^{p-1}\left(|a|^{p}+|b|^{p}\right)$ for all $a,b\in\mathbb{R}$ and $p>1$ yields:
$\left(\frac{\widetilde{T}^{S}}{P^{S}}\right)^{4}$ $=\left(\frac{\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\cdot\left|\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\mid\left(S_{i,g,t}^{\ast}\right)\right]\right|}{\frac{1}{G}\sum_{(i,g,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}}\right)^{4}$
$\leq\left(\max_{(i,g,t):t\geq 2,1\leq g\leq G}\left|\mathrm{E}\left[Y_{i,g,t}(1)-Y_{i,g,t}(0)\mid\left(S_{i,g,t}^{\ast}\right)\right]\right|\right)^{4}$
$\leq\max_{(i,g,t):t\geq 2,1\leq g\leq G}\left(\mathrm{E}\left[\left|Y_{i,g,t}(1)-Y_{i,g,t}(0)\right|^{4}\mid\left(S_{i,g,t}^{\ast}\right)\right]\right)$
$\leq 16\sup_{(i,g,t):t\geq 2,g\in\mathbb{N}\setminus\{0\}}\mathrm{E}\left[Y_{i,g,t}(d)^{4}\mid\left(S_{i,g,t}^{\ast}\right)\right]$
$\leq C_{3},$
for some constant $C_{3}>0$ by Assumption 18. Moreover, since the above bound holds uniformly in $G$, it holds that
$\sup_{G\in\mathbb{N}\setminus\{0\}}\mathrm{E}\left[\left(\frac{\widetilde{T}^{S}}{P^{S}}\right)^{4}\right]\leq C_{3}.$
On the other hand, we observe that the random variables $\left\{\sum_{(i,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\right\}$ are independent across $g$ under Assumption 14. Moreover, by Assumption 18, for each $g\in\{1,\dots,G\}$, there exists constant $C_{4}>0$ such that
$0\leq\sum_{(i,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\leq\sum_{t\geq 2}N_{g,t}\leq C_{4}.$
Hence, applying Hoeffding’s inequality for the sum of bounded independent random variables yields:
$\mathrm{Pr}(I_{G}=0)$ $=\mathrm{Pr}\left(\left|P^{S}-\mathrm{E}\left[P^{S}\right]\right|\geq\varepsilon_{G}\right)$
$=\mathrm{Pr}\left(\left|\sum_{g=1}^{G}\sum_{(i,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}-\mathrm{E}\left[\sum_{g=1}^{G}\sum_{(i,t):t\geq 2}\mathbb{I}\{S_{i,g,t}^{*}=1\}\right]\right|\geq G\varepsilon_{G}\right)$
$\leq 2\exp\left(-\frac{2G^{2}\varepsilon_{G}^{2}}{\sum_{g=1}^{G}C_{4}^{2}}\right)=2\exp\left(-\frac{2G\varepsilon_{G}^{2}}{C_{4}^{2}}\right).$
Now, choose $\varepsilon_{G}:=\sqrt{\frac{\log G}{G}}$. Then,
$\mathrm{Pr}(I_{G}=0)\leq 2G^{-2/C_{4}^{2}},$
which implies
$\lim_{G\rightarrow\infty}B_{2}=\lim_{G\rightarrow\infty}\left(\mathrm{E}\left[\left(\frac{\widetilde{T}^{S}}{P^{S}}\right)^{4}\right]\cdot\mathrm{Pr}(I_{G}=0)\right)^{1/4}\leq\lim_{G\rightarrow\infty}\left(C_{3}\cdot 2G^{-2/C_{4}^{2}}\right)^{1/4}=0.$
Hence,
$\sqrt{G}\cdot B=o(1).$
Finally, consider the first term $A$ on the right-hand side of (26). We decompose:
$A$ $=\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{\widetilde{T}^{S}}{P^{S}}\cdot I_{G}\right]$
$=\underbrace{\mathrm{E}\left[\left(\widetilde{T}^{S}-\mathrm{E}\left[\widetilde{T}^{S}\right]\right)\cdot\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{I_{G}}{P^{S}}\right]}_{=:A_{1}}+\underbrace{\mathrm{E}\left[\widetilde{T}^{S}\right]\cdot\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{I_{G}}{P^{S}}\right]}_{=:A_{2}}.$
For the term $A_{1}$, we apply Cauchy-Schwarz:
$\left|A_{1}\right|$ $=\left|\mathrm{E}\left[\left(\widetilde{T}^{S}-\mathrm{E}\left[\widetilde{T}^{S}\right]\right)\cdot\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{I_{G}}{P^{S}}\right]\right|$
$\leq\left(\mathrm{Var}\left[\widetilde{T}^{S}\right]\right)^{1/2}\cdot\left(\mathrm{E}\left[\left(\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{I_{G}}{P^{S}}\right)^{2}\right]\right)^{1/2}.$
Observe that on the event $\left\{I_{G}=1\right\}$, we have $\left|P^{S}-\mathrm{E}\left[P^{S}\right]\right|<\varepsilon_{G}$ with $\varepsilon_{G}=\sqrt{\frac{\log G}{G}}=o(1)$. This implies:
$P^{S}>\mathrm{E}\left[P^{S}\right]-\varepsilon_{G},$
where $\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}>0$ for sufficiently large $G$. Consequently,
$\left(\mathrm{E}\left[\left(\frac{P^{S}-\mathrm{E}\left[P^{S}\right]}{P^{S}}\cdot I_{G}\right)^{2}\right]\right)^{1/2}\leq\left(\mathrm{E}\left[\left(\frac{P^{S}-\mathrm{E}\left[P^{S}\right]}{P^{S}}\right)^{2}\right]\right)^{1/2}\leq\left(\mathrm{E}\left[\left(\frac{P^{S}-\mathrm{E}\left[P^{S}\right]}{\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}}\right)^{2}\right]\right)^{1/2}\leq\frac{\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}{\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}}.$
By Assumption 18, the quantities $\mathrm{E}\left[P^{S}\right]$ , $\sqrt{G}\cdot\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}$ and
$\sqrt{G}\cdot\left(\mathrm{Var}\left[\widetilde{T}^{S}\right]\right)^{1/2}$ $=\sqrt{G}\cdot\left(\mathrm{Var}\left[\mathrm{E}\left[T^{S}\mid\left(S_{i,g,t}^{\ast}\right)\right]\right]\right)^{1/2}$
$=\sqrt{G}\cdot\left(\mathrm{Var}\left[T^{S}\right]-\mathrm{E}\left[\mathrm{Var}\left[T^{S}\mid\left(S_{i,g,t}^{\ast}\right)\right]\right]\right)^{1/2}$
$\leq\sqrt{G}\cdot\left(\mathrm{Var}\left[T^{S}\right]\right)^{1/2}$
converge to finite limits. Thus, for sufficiently large $G$,
$\sqrt{G}\cdot|A_{1}|\leq\sqrt{G}\cdot\left(\mathrm{Var}\left[T^{S}\right]\right)^{1/2}\cdot\frac{\sqrt{G}\cdot\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}{\sqrt{G}\cdot\left(\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}\right)}\xrightarrow{G\rightarrow\infty}0.$
It remains to examine the term $A_{2}$. We rewrite:
$A_{2}=\mathrm{E}\left[\widetilde{T}^{S}\right]\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\frac{I_{G}}{P^{S}}\right]=-\mathrm{E}\left[\widetilde{T}^{S}\right]\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\frac{1-I_{G}}{P^{S}}\right].$
Recalling what we have derived above:
$\mathrm{Pr}(I_{G}=0)\leq 2G^{-2/C_{4}^{2}}\quad\text{and}\quad\left(\mathrm{E}\left[\left(\frac{P^{S}-\mathrm{E}\left[P^{S}\right]}{P^{S}}\right)^{2}\right]\right)^{1/2}\leq\frac{\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}{\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}}$
for sufficiently large $G$. Thus, applying Cauchy–Schwarz yields:
$\left|A_{2}\right|$ $=\left|\mathrm{E}\left[\widetilde{T}^{S}\right]\cdot\mathrm{E}\left[\left(P^{S}-\mathrm{E}\left[P^{S}\right]\right)\cdot\frac{1-I_{G}}{P^{S}}\right]\right|$
$\leq\left|\mathrm{E}\left[\widetilde{T}^{S}\right]\right|\cdot\left(\mathrm{E}\left[\left(\frac{P^{S}-\mathrm{E}\left[P^{S}\right]}{P^{S}}\right)^{2}\right]\right)^{1/2}\cdot\left(\mathrm{E}\left[\left(1-I_{G}\right)^{2}\right]\right)^{1/2}$
$\leq\left|\mathrm{E}\left[\widetilde{T}^{S}\right]\right|\cdot\frac{\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}{\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}}\cdot 2G^{-1/C_{4}^{2}}.$
Moreover, $\mathrm{E}\left[\widetilde{T}^{S}\right]=\mathrm{E}\left[T^{S}\right]$ by definition, and the quantities $\mathrm{E}\left[P^{S}\right]$ , $\sqrt{G}\cdot\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}$ and $\mathrm{E}\left[T^{S}\right]$ converges to a finite limit by Assumption 18. For sufficiently large $G$,
$\sqrt{G}\cdot|A_{2}|\leq\left|\mathrm{E}\left[T^{S}\right]\right|\cdot\frac{\sqrt{G}\cdot\left(\mathrm{Var}\left[P^{S}\right]\right)^{1/2}}{\mathrm{E}\left[P^{S}\right]-\varepsilon_{G}}\cdot 2G^{-1/C_{4}^{2}}\xrightarrow{G\rightarrow\infty}0.$
This completes the proof. ∎
Proof of Proposition 7.
(i) We only examine the asymptotic behavior of $T_{1,0}^{s}$ under $H_{0}^{MC}$. The arguments for $T_{1,0}^{m}$ are analogous. To examine the asymptotic behavior of $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$ under $H_{0}^{MC}$ defined in (18), recall the definition of $\boldsymbol{U}$ and note that for some generic constants $C>0$ it holds that:
$\sup_{g,t}N_{g,t}^{4}\mathrm{E}\left[1\{D_{g,t}=1,D_{g,t-1}=0\}(Y_{g,t}-Y_{g,t-l})^{4}\right]$
$\leq$ $\sup_{g,t}N_{g,t}^{4}\times\sup_{g,t}\mathrm{E}\left[(Y_{g,t}-Y_{g,t-l})^{4}\right]$
$\leq$ $C\sup_{g,t}N_{g,t}^{4}\times\sup_{g,t}\mathrm{E}\left[Y_{g,t}^{4}\right]<+\infty,$
where the last inequality follows from Assumption 18. Thus, by another application of Assumption 18 and the Lypuanov’s Central Limit Theorem:
$\sqrt{G}\left(\mathbf{U}-\mathrm{E}\left[\mathbf{U}\right]\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{U}}),$
where $\boldsymbol{\Sigma}_{\mathbf{U}}$ is defined in (22). Next, define:
$\boldsymbol{f}\left(\begin{array}[]{c}p_{0,0}(t,l),q_{0,0}(t,l),p_{1,0}(t,l),q_{1,0}(t,l)\\ \ldots\\ p_{0,0}(j,k),q_{0,0}(j,k),p_{1,0}(j,k),q_{1,0}(j,k)\end{array}\right)\equiv\left(\begin{array}[]{c}\frac{q_{1,0}(t,l)-\frac{p_{1,0}(t,l)}{p_{0,0}(t,l)}q_{0,0}(t,l)}{p_{1,0}(t,l)}\\ \ldots\\ \frac{q_{1,0}(j,k)-\frac{p_{1,0}(j,k)}{p_{0,0}(j,k)}q_{0,0}(j,k)}{p_{1,0}(j,k)}\end{array}\right)$
for all elements of $\mathcal{C}_{S}$ such that the denominator is strictly positive. Moreover, let
$\boldsymbol{f}\left(\mathrm{E}\left[\mathbf{U}\right]\right)=\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D},$
where
$\boldsymbol{f}_{N}=\left(\begin{array}[]{c}f_{N}(t,l)\\ \ldots\\ f_{N}(j,k)\end{array}\right)$
with elements:
$f_{N}(t,l)=$ $\frac{1}{G}\sum_{g\in G}N_{g,t}\left(\mathrm{E}\left[1\{D_{g,t}=1,D_{g,t-1}=0\}(Y_{g,t}-Y_{g,t-l})\right]\right.$
$\left.-\frac{\mathrm{E}\left[P_{1,0}(t,l)\right]}{\mathrm{E}\left[P_{0,0}(t,l)\right]}\mathrm{E}\left[1\{D_{g,t}=1,D_{g,t-1}=0\}(Y_{g,t}-Y_{g,t-l})\right]\right)$
and
$\boldsymbol{f}_{D}=\left(\begin{array}[]{c}f_{D}(t,l)\\ \ldots\\ f_{D}(j,k)\end{array}\right)=\left(\begin{array}[]{c}\frac{1}{\mathrm{E}\left[P_{1,0}(t,l)\right]}\\ \ldots\\ \frac{1}{\mathrm{E}\left[P_{1,0}(j,k)\right]}\end{array}\right)^{\prime}.$
First, note that under $H_{0}^{MC}$:
$\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$
$=$ $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{\tau}_{1,0}\right)$
$=$ $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D}\right)+\sqrt{G}\left(\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D}-\boldsymbol{\tau}_{1,0}\right),$
where $\boldsymbol{\tau}_{1,0}$ was defined in (19). By Assumption 19, we have that $\mathrm{E}\left[\mathbf{U}(t,l)\right]$ converges pointwise for every $t$ and $l$ to $\mathbf{U}^{\infty}(t,l)$. Thus, also $\mathrm{E}\left[\mathbf{U}\right]$ converges to $\mathbf{U}^{\infty}$. Moreover, as $\boldsymbol{f}$ is differentiable in a neighborhood of $\mathbf{U}^{\infty}$, it follows again by the uniform Delta method (e.g., Theorem 3.8 Van der Vaart, 1998) that:
$\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D}\right)=\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)\times\sqrt{G}\left(\mathbf{U}-\mathrm{E}\left[\mathbf{U}\right]\right)+o_{\Pr}(1),$
where:
$\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)=\left(\begin{array}[]{ccccc}\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(t,l)\right)^{\prime}&\mathbf{0}_{4}&\ldots&\ldots&\mathbf{0}_{4}\\ \mathbf{0}_{4}&\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(t,m)\right)^{\prime}&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(m,k)\right)^{\prime}&\mathbf{0}_{4}\\ \mathbf{0}_{4}&\ldots&\ldots&\mathbf{0}_{4}&\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(j,k)\right)^{\prime}\end{array}\right)$ (27)
is a $|\mathcal{C}_{S}|\times(4\cdot|\mathcal{C}_{S}|)$ matrix each element $\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(t,l)\right)$ is given by:
$\mathbf{J}_{f}\left(\mathbf{U}^{\infty}(t,l)\right)=\lim_{G\rightarrow\infty}\left(\begin{array}[]{c}\frac{\mathrm{E}\left[Q_{0,0}(t,l)\right]}{\mathrm{E}\left[P_{0,0}(t,l)^{2}\right]}\\ -\frac{1}{\mathrm{E}\left[P_{0,0}(t,l)\right]}\\ -\frac{\mathrm{E}\left[Q_{1,0}(t,l)\right]}{\mathrm{E}\left[P_{1,0}(t,l)^{2}\right]}\\ \frac{1}{\mathrm{E}\left[P_{1,0}(t,l)\right]}\end{array}\right)$
and $\mathbf{0}_{4}$ denotes a 1×41$\times$ 41 × 4 vector of zeros. We thus have that:
$\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{f}\left(\mathrm{E}\left[\mathbf{U}\right]\right)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\mathbf{0},\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)\boldsymbol{\Sigma}_{\mathbf{U}}\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)^{\prime}\right),$
[…]
$\sqrt{G}\left(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,\sigma^{2}(t,l)\right),$
where:
$\sigma^{2}(t,l):=\Omega_{1,0}^{(1)}(t,l)+\Omega_{1,0}^{(2)}(t,l)+2\Omega_{1,0}^{(12)}(t,l)+\Omega_{0,0}^{(1)}(t,l)+\Omega_{0,0}^{(2)}(t,l)+2\Omega_{0,0}^{(12)}(t,l),$ (28)
with:
$\Omega_{1,0}^{(1)}(t,l)=\left(-\frac{\mathrm{E}_{\infty}\left[Q_{1,0}(t,l)\right]}{\mathrm{E}_{\infty}\left[P_{1,0}(t,l)^{2}\right]}\right)^{2}\text{Var}_{\infty}\left[P_{1,0}(t,l)\right],$
$\Omega_{1,0}^{(2)}(t,l)=\left(\frac{1}{\mathrm{E}_{\infty}\left[P_{1,0}(t,l)^{2}\right]}\right)^{2}\text{Var}_{\infty}\left[Q_{1,0}(t,l)\right],$
and
$\Omega_{1,0}^{(12)}(t,l)=\left(-\frac{\mathrm{E}_{\infty}\left[Q_{1,0}(t,l)\right]}{\mathrm{E}_{\infty}\left[P_{1,0}(t,l)^{2}\right]\mathrm{E}_{\infty}\left[P_{1,0}(t,l)\right]}\right)\text{Cov}_{\infty}\left[Q_{1,0}(t,l),P_{1,0}(t,l)\right]$
(the terms $\Omega_{0,0}^{(1)}(t,l)$, $\Omega_{0,0}^{(2)}(t,l)$, $\Omega_{0,0}^{(12)}(t,l)$ are defined analogously). Replicating the arguments from the proof of Proposition 6, we also obtain that:
$\sqrt{G}\left(\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D}-\boldsymbol{\tau}_{1,0}\right)=\sqrt{G}\left(\boldsymbol{f}_{N}\cdot\boldsymbol{f}_{D}-\mathbf{0}\right)=o_{\Pr}(1).$
Next, we show the first order asymptotic validity of the bootstrap moment equalities under the null hypothesis. Specifically, we want to show that under $H_{0}^{MC}$, for any $\nu>0$:
$\Pr\left(\sup_{\mathbf{x}\in\mathbb{R}^{\mathcal{C}_{S}}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\boldsymbol{\tau}}_{1,0}^{b}-\widehat{\boldsymbol{\tau}}_{1,0})\leq\mathbf{x}\right)-\Pr\left(\sqrt{G}\widehat{\boldsymbol{\tau}}_{1,0}\leq\mathbf{x}\right)\right|>\nu\right)\rightarrow 0,$
where the inequality is understood elementwise. We will do so by showing that for each element $(t,l)$ of $\mathcal{C}_{S}$ pointwise:
$\Pr\left(\sup_{x\in\mathbb{R}}\left|\operatorname*{\text{Pr}^{b}}\left(\sqrt{G}(\widehat{\tau}_{1,0}^{b}(t,l)-\widehat{\tau}_{1,0}(t,l))\leq x\right)-\Pr\left(\sqrt{G}\widehat{\tau}_{1,0}(t,l)\leq x\right)\right|>\nu\right)\rightarrow 0,$ (29)
Since the cardinality of $\mathcal{C}_{S}$ is finite, the uniform result will follow by the Cramér-Wold device once we establish that that the bootstrap covariances adequately mimic the actual covariances of the moment equalities.
We start with the convergence in (29). To show this result, we establish that for each element $(t,l)$ in $\mathcal{C}_{S}$:
$\sqrt{G}(\widehat{\tau}_{1,0}^{b}(t,l)-\widehat{\tau}_{1,0}(t,l))\stackrel{{\scriptstyle d^{b}}}{{\rightarrow}}N\left(0,\sigma^{2}(t,l)\right),$
in probability. Then, since:
$\sqrt{G}\left(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,\sigma^{2}(t,l)\right),$
the result will follow from the Triangle inequality and Polya’s Theorem given the continuity of the normal distribution. As in the first part, we start by arguing that:
$\sqrt{G}\left(\mathbf{U}^{b}(t,l)-\mathrm{E}^{b}\left[\mathbf{U}^{b}(t,l)\right]\right)\stackrel{{\scriptstyle d^{b}}}{{\rightarrow}}N(\mathbf{0},\boldsymbol{\Sigma}(t,l)),$
in probabilty, which we achieve by applying Lyapunov’s CLT for triangular arrays as the bootstrap observations are row-wise independent. To establish the latter result, first note that as the bootstrap mean equals the sample mean:
$\mathrm{E}^{b}\left[Q_{j,0}^{b}(t,l)\right]=Q_{1,0}(t,l)$
and
$\mathrm{E}^{b}\left[P_{1,0}^{b}(t,l)\right]=P_{1,0}(t,l),$
so that:
$\mathrm{E}^{b}\left[Q_{j,0}^{b}(t,l)\right]-Q_{j,0}(t,l)=0$
as well as
$\mathrm{E}^{b}\left[P_{j,0}^{b}(t,l)\right]-P_{j,0}(t,l)=0$
for $j\in\{0,1\}$. Secondly, focusing on $Q_{1,0}^{b}(t,l)$ for illustrative purposes, because the (block) bootstrap observations are i.i.d. across groups, conditional on the sample, we have that:
$\text{Var}^{b}\left[Q_{1,0}^{b}(t,l)\right]$
$=$ $\frac{1}{G}\sum_{g\in G}\left\{N_{g,t}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\left\{\left(Y_{g,t-1}-Y_{g,t-l}\right)\right.\right.$
$\left.\left.-\left(\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right)\right)\right\}\right\}^{2}$
$\stackrel{{\scriptstyle\Pr}}{{\rightarrow}}$ $\mathrm{Var}_{\infty}\left[Q_{1,0}(t,l)\right],$
and the same argument applies to the other expressions as well. Focusing again on $Q_{1,0}^{b}(t,l)$ for simplicity and denoting $Z_{g}^{b}:=\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)$, it remains to analyze the Lyapunov condition:
$\frac{1}{\sum_{g\in G}\mathrm{E}^{b}\left[\left|Z_{g}^{b}\right|^{2}\right]^{1+\delta/2}}\sum_{g\in G}\mathrm{E}^{b}\left[\left|Z_{g}^{b}\right|^{2+\delta}\right]\stackrel{{\scriptstyle\mathrm{Pr}}}{{\rightarrow}}0$
for some $\delta>0$. That is:
$\frac{1}{\left(\sum_{g\in G}\left|Z_{g}\right|^{2}\right)^{1+\delta/2}}\sum_{g\in G}\left(\left|Z_{g}\right|^{2+\delta}\right)\stackrel{{\scriptstyle\mathrm{Pr}}}{{\rightarrow}}0.$
Thus, first define $\overline{Z}_{g}=|Z_{g}|^{2+\delta}-\mathrm{E}\left[|Z_{g}^{b}|^{2+\delta}\right]$, and note that for some $\epsilon>0$:
$\mathrm{Pr}\left[\frac{1}{G}\sum_{g}\left|\overline{Z}_{g}\right|<\epsilon\right]$
$\leq$ $\frac{\mathrm{E}\left[\left(\sum_{g}\overline{Z}_{g}\right)^{4}\right]}{G^{4}\epsilon^{4}}$
$\leq$ $\frac{C\left(\sum_{g}\mathrm{E}\left[\overline{Z}_{g}^{4}\right]+\left(\sum_{g}\mathrm{E}\left[\overline{Z}_{g}^{2}\right]\right)^{2}\right)}{G^{4}\epsilon^{4}},$
where the last line follows from Rosenthal’s inequality for some $C>1$. The RHS of the last line is bounded, however, by the moment conditions in Assumption 19 since:
$\mathrm{E}\left[\overline{Z}_{g}^{4}\right]$
$=$ $\mathrm{E}\left[\left(|Z_{g}|^{2+\delta}-\mathrm{E}\left[|Z_{g}^{b}|^{2+\delta}\right]\right)^{4}\right]$
$=$ $O\left(\mathrm{E}\left[|Z_{g}|^{8+4\delta}\right]\right)$
$\leq$ $O\left(\mathrm{E}\left[\sup_{g,t}|Y_{g,t}|^{8+4\delta}\right]\right)$
$=$ $O(1).$
This in turn implies that
$\sum_{g\in G}\mathrm{E}^{b}\left[\left|Z_{g}^{b}\right|^{2+\delta}\right]=O_{\mathrm{Pr}}\left(G\right).$
Similarly, for the denominator we have that:
$\sum_{g\in G}\mathrm{E}^{b}\left[\left|Z_{g}^{b}\right|^{2}\right]^{1+\delta/2}=O_{\mathrm{Pr}}\left(G^{1-\delta/2}\right),$
which implies that Lyapunov’s condition is satisfied. It follows that:
$\sqrt{G}\left(Q_{1,0}^{b}(t,l)-\mathrm{E}^{b}\left[Q_{1,0}^{b}(t,l)\right]\right)\stackrel{{\scriptstyle d^{b}}}{{\rightarrow}}N(\mathbf{0},\mathrm{Var}_{\infty}\left[Q_{1,0}(t,l)\right]),$
in probability, and since the same argument can also be made for the other terms that:
$\sqrt{G}\left(\mathbf{U}^{b}(t,l)-\mathrm{E}^{b}\left[\mathbf{U}^{b}(t,l)\right]\right)\stackrel{{\scriptstyle d^{b}}}{{\rightarrow}}N(\mathbf{0},\boldsymbol{\Sigma}(t,l)),$
in probability. Moreover, since $f(t,l)=f_{N}(t,l)\cdot f_{D}(t,l)$ is differentiable, by the uniform Delta method and the same arguments as in the first part of the proof:
$\sqrt{G}\left(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,\sigma^{2}(t,l)\right),$
in probability, with $\sigma^{2}(t,l)$ defined in (28).
It remains to analyze the covariance of different elements of the vector $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}^{b}-\widehat{\boldsymbol{\tau}}\right)$. For simplicity, we only examine the terms involving (1,0)(1,0)( 1 , 0 ):
$\text{Cov}^{b}\left[\sqrt{G}\left(\frac{Q_{1,0}^{b}(t,l)}{P^{b}_{1,0}(t,l)}-\frac{Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right),\sqrt{G}\left(\frac{Q_{1,0}^{b}(j,k)}{P^{b}_{1,0}(j,k)}-\frac{Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right]$
$=$ $\mathrm{E}^{b}\left[G\left(\frac{Q_{1,0}^{b}(t,l)}{P^{b}_{1,0}(t,l)}-\frac{Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right)\left(\frac{Q_{1,0}^{b}(j,k)}{P^{b}_{1,0}(j,k)}-\frac{Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right]$
$-\mathrm{E}^{b}\left[\sqrt{G}\left(\frac{Q_{1,0}^{b}(t,l)}{P^{b}_{1,0}(t,l)}-\frac{Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right)\right]\times\mathrm{E}^{b}\left[\sqrt{G}\left(\frac{Q_{1,0}^{b}(j,k)}{P^{b}_{1,0}(j,k)}-\frac{Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right]$
$=$ $\mathrm{E}^{b}\left[G\left(\frac{Q_{1,0}^{b}(t,l)}{P^{b}_{1,0}(t,l)}-\frac{Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right)\left(\frac{Q_{1,0}^{b}(j,k)}{P^{b}_{1,0}(j,k)}-\frac{Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right],$
where the equality follows since, conditional on the sample, the bootstrap mean of the recentered terms is zero. Straightforward algebra also shows that the above term can be rewritten as:
$\mathrm{E}^{b}\left[G\left(\frac{Q_{1,0}^{b}(t,l)-Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right)\left(\frac{Q_{1,0}^{b}(j,k)-Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right]$
$+$ $\mathrm{E}^{b}\left[G\left(\frac{1}{P^{b}_{1,0}(t,l)}-\frac{1}{P_{1,0}(t,l)}\right)\left(\frac{1}{P^{b}_{1,0}(j,k)}-\frac{1}{P_{1,0}(j,k)}\right)Q_{1,0}^{b}(t,l)Q_{1,0}^{b}(j,k)\right]$
$=$ $I_{1,0}^{b}+II_{1,0}^{b}$
We will only examine $I_{1,0}^{b}$ in what follows, the term $II_{1,0}^{b}$ follows by similar arguments. Moreover, let
$m(t,l)\equiv\frac{1}{G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right),$
noting that $m(t,l)=\mathrm{E}^{b}\left[\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)\right]$. Thus, observe that:
$\mathrm{E}^{b}\left[G\left(\frac{Q_{1,0}^{b}(t,l)-Q_{1,0}(t,l)}{P_{1,0}(t,l)}\right)\left(\frac{Q_{1,0}^{b}(j,k)-Q_{1,0}(j,k)}{P_{1,0}(j,k)}\right)\right]$
$=$ $\mathrm{E}^{b}\left[\frac{1}{P_{1,0}(t,l)\sqrt{G}}\sum_{g\in G}\left(\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-m(t,l)\right)\right.$
×$\left.\times\frac{1}{P_{1,0}(j,k)\sqrt{G}}\sum_{g\in G}\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
$=$ $\mathrm{E}^{b}\left[\frac{1}{P_{1,0}(t,l)P_{1,0}(j,k)G}\sum_{g\in G}\left(\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-m(t,l)\right)\right.$
×$\left.\times\left(\mathbb{I}\{D^{b}_{g,j}=1,D^{b}_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}^{b}-Y_{g,j-k}^{b}\right)-m(j,k)\right)\right]$
$+\mathrm{E}^{b}\left[\frac{1}{P_{1,0}(t,l)P_{1,0}(j,k)G}\sum_{g\in G}\sum_{h\neq g}\left(\mathbb{I}\{D^{b}_{g,t}=1,D^{b}_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}^{b}-Y_{g,t-l}^{b}\right)-m(t,l)\right)\right.$
×$\left.\times\left(\mathbb{I}\{D^{b}_{h,j}=1,D^{b}_{h,j-1}=0\}N_{h,j}\left(Y_{h,j-1}^{b}-Y_{h,j-k}^{b}\right)-m(j,k)\right)\right]$
$=$ $\frac{1}{P_{1,0}(t,l)P_{1,0}(j,k)G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}\mathbb{I}\{D_{g,j}=1,D_{g,j-1}=0\}N_{g,t}N_{g,j}$
×$\times\left(Y_{g,t-1}-Y_{g,t-l}\right)\left(Y_{g,j-1}-Y_{g,j-k}\right)$
$-\left(\frac{1}{P_{1,0}(t,l)G}\sum_{g\in G}\mathbb{I}\{D_{g,t}=1,D_{g,t-1}=0\}N_{g,t}\left(Y_{g,t-1}-Y_{g,t-l}\right)\right)$
×$\times\left(\frac{1}{P_{1,0}(j,k)G}\sum_{g\in G}\mathbb{I}\{D_{g,j}=1,D_{g,j-1}=0\}N_{g,j}\left(Y_{g,j-1}-Y_{g,j-k}\right)\right),$
where the last equality follows since the blocks are, conditional on the sample, independent. We thus have that the bootstrap covariance is equal to the sample covariance, which converges in probability to $\lim_{G\rightarrow\infty}\mathrm{Cov}\left[\sqrt{G}(\widehat{\tau}_{1,0}(t,l)-\tau_{1,0}(t,l)),\sqrt{G}(\widehat{\tau}_{1,0}(j,k)-\tau_{1,0}(j,k))\right]$. Uniform convergence now follows by an application of the Cramér-Wold device, see for instance White (2000). Proposition 7 (i) then follows from continuous mapping.
(ii) Under the alternative $H_{A}^{MC}$:
$\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\mathbf{0}\right)$
$=$ $\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{\tau}_{1,0}\right)+\sqrt{G}\left(\boldsymbol{\tau}_{1,0}-\mathbf{0}\right)$
For the first term on the RHS, it follows by the arguments of part (i) that:
$\sqrt{G}\left(\widehat{\boldsymbol{\tau}}_{1,0}-\boldsymbol{\tau}_{1,0}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\mathbf{0},\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)\boldsymbol{\Sigma}_{\mathbf{U}}\mathbf{J}_{f}\left(\mathbf{U}^{\infty}\right)^{\prime}\right).$
On the other hand, for the second term, we have that:
$\left|\sqrt{G}\left(\boldsymbol{\tau}_{1,0}-\mathbf{0}\right)\right|\rightarrow+\infty.$
Since the bootstrap mimics the distribution under $H_{0}^{MC}$, the result follows. ∎